### v1

# Yongtao Li, Lihua Feng, Yuejian Peng

Abstract
A fundamental result in extremal graph theory attributes to Mantel’s theorem, which states that every graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. About half of a century ago, Lovász and Simonovits (1975) provided a supersaturation phenomenon, which asserts that for $q<n/2$, every graph with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles. This result solved a conjecture proposed by Erdős in 1962. In this paper, we establish a spectral version of the result of Lovász and Simonovits. Let $Y_{n,2,q}$ be the graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding a matching with $q$ edges into the vertex part of size $\lceil n/2\rceil$. Using the supersaturation-stability method and the classical spectral techniques, we firstly prove that for $n\geq 300q^{2}$, each graph $G$ on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$ contains at least $q\lfloor n/2\rfloor$ triangles. Moreover, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star with $q$ edges into the vertex part of size $\lceil n/2\rceil$. Secondly, we show further that $T_{n,2,q}$ is the unique spectral extremal graph that contains at most $q\lfloor n/2\rfloor$ triangles and attains the maximum of the spectral radius. This result answers a spectral triangle counting problem due to Ning and Zhai (2023). Thirdly, we present an asymptotically spectral stability result under a specific constraint on the triangle covering number. The third result could be regarded as a spectral extension of a recent result proved by Balogh and Clemen (2023), and independently by Liu and Mubayi (2022).
Key words: Extremal graph theory; counting triangles; adjacency matrix.
2010 Mathematics Subject Classification. 05C35; 05C50.
1 Introduction
The Turán type problem states that for a graph $F$, what is the minimum integer $m$ such that every $n$-vertex graph with more than $m$ edges contains a copy of $F$. This extremal number is denoted by $\mathrm{ex}(n,F)$. It is a longstanding open problem in extremal combinatorics to develop some understanding on the function $\mathrm{ex}(n,F)$ for a general graph $F$. A well-known result of Mantel (see [4, 33]) asserts that an $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. The bound is tight by considering the bipartite Turán graph $T_{n,2}$, which is a complete bipartite graph whose two parts have sizes as equal as possible. Shortly afterwards, there have been various generalizations of Mantel’s theorem; see, e.g., [5, 6].
1.1 Extremal graph problems on triangles
About seventy years ago, Rademacher (unpublished, see Erdős [15, 18]) showed an interesting extension on Mantel’s theorem. Namely, there are not just one but at least $\lfloor{n}/{2}\rfloor$ triangles in any $n$-vertex graph with more than $\lfloor{n^{2}}/{4}\rfloor$ edges.
Theorem 1.1 (Erdős–Rademacher, 1955).
If $G$ is a graph on $n$ vertices with
$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+1,$
then $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G$ is the graph obtained from $T_{n,2}$ by adding one edge into the vertex part of size $\lceil{n}/{2}\rceil$.
We remark that Erdős [18] proved a variant for a graph with less than $\lfloor n^{2}/4\rfloor$ edges. The result of Erdős states that if $\ell\geq 0$ and $G$ is an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor-\ell$, and $G$ has a triangle, then it has at least $\lfloor{n}/{2}\rfloor-\ell-1$ triangles. As a special case, we can see that if $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$, then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G=T_{n,2}$. The problem of determining the minimum number of triangles is usually referred to as the Erdős–Rademacher problem, which is regarded as the starting point of the supersaturation problem in extremal graph theory. After proving Theorem 1.1, Erdős [16, 17] showed that there exists a small constant $c>0$ such that if $1\leq q<cn$ and $n$ is sufficiently large, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles. This result was further generalized by Lovász and Simonovits [34, 35] as follows.
Theorem 1.2 (Lovász–Simonovits, 1975).
Let $1\leq q<{n}/{2}$ be an integer and $G$ be an $n$-vertex graph with
$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+q.$
Then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles.
A quick proof of Theorem 1.2 in the range of $q\leq{n}/{4}$ can be found in [4, page 302]. Incidentally, we mention that the condition $q<{n}/{2}$ is necessary. For even $n$ and $q={n}/{2}$, one can add $q+1$ extra edges to the larger vertex part of $K_{\frac{n}{2}+1,\frac{n}{2}-1}$ such that there is no triangle among these $q+1$ edges. Hence, we get a graph with ${n^{2}}/{4}+q$ edges and $(q+1)({n}/{2}-1)={n^{2}}/{4}-1$ triangles, which is less than $q\lfloor{n}/{2}\rfloor$ triangles.
The stabilities of Theorems 1.1 and 1.2 were recently studied by Xiao and Katona [56], Liu and Mubayi [32] and Balogh and Clemen [3]. A graph is called color-critical if it has an edge whose deletion reduces its chromatic number. For example, the complete graph, odd cycle, book and even wheel are color-critical. Apart from the triangles, the supersaturation problem for color-critical graphs was also investigated by Mubayi [37], Pikhurko and Yilma [48], Ma and Yuan [36]. For related results, we refer to [49, 30] and references therein.
1.2 Eigenvalues and counting triangles
Spectral graph theory aims to apply the eigenvalues of associated matrices to investigate the structural information of a graph. Let $G$ be a simple graph on $[n]$. The adjacency matrix is defined as $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $i$ and $j$ are adjacent, and $a_{i,j}=0$ otherwise. Let $\lambda(G)$ be the spectral radius of $G$, which is defined as the maximum of modulus of eigenvalues of $A(G)$. By the Perron–Frobenius theorem, we know that $\lambda(G)$ is the largest eigenvalue of $A(G)$ and there exists a nonnegative eigenvector corresponding to $\lambda(G)$.
As mentioned before, the Turán type problem studies the condition on the number of edges of a graph to force certain subgraphs, e.g., the classical Turán theorem states that $e(G)>(1-\nicefrac{{1}}{{r}})n^{2}/2$ forces a clique $K_{r+1}$. Correspondingly, one could wish to investigate the condition on spectral radius of the associated adjacency matrix to guarantee certain subgraphs. The interplay between these two areas above is called the spectral Turán type problem. Finding general conditions on the spectral radius which ensure that $G$ has specific substructures or properties is one of the most popular topic in Spectral Extremal Graph Theory. One of the well-known results on this topic was obtained in 1986 by Wilf [54] who showed that every $n$-vertex graph $G$ with $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$ has a clique $K_{r+1}$. Another significant result due to Nikiforov [38] states that an $m$-edge graph $G$ with $\lambda^{2}(G)>{(1-\nicefrac{{1}}{{r}})2m}$ contains a clique $K_{r+1}$. The above spectral results generalize the Turán theorem. As illustrated by Nikiforov, his result can imply Wilf’s result by assuming Turán’s theorem. Actually, this implication can also be seen by the fact $\lambda(G)\geq 2m/n$. Indeed, we know that if $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$, then $\lambda^{2}(G)>(1-\nicefrac{{1}}{{r}})n\cdot(2m/n)=(1-\nicefrac{{1}}{{r}})2m$, as expected. In addition, the spectral radius of a graph is always at least the minimum degree, if $G$ is a $K_{r+1}$-free graph on $n$ vertices, then Wilf’s result implies $\delta(G)\leq(1-\nicefrac{{1}}{{r}})n$ and so $\delta(G)\leq n-\lceil n/r\rceil=\delta(T_{n,r})$. Taking a vertex $u\in V(G)$ with degree $d(u)=\delta(G)$, it follows by induction that $e(G)=e(G\setminus\{u\})+\delta(G)\leq e(T_{n-1,r})+\delta(T_{n,r})=e(T_{n,r})$. We conclude that Wilf’s result can imply the “strong” Turán theorem as well.
The spectral Turán problems have been receiving considerable attention in the last two decades and it is an important topic in modern graph theory; see, e.g., [38, 7, 23] for cliques, [20] for matchings, [12] for even cycles, [21, 62, 22] for cycles of consecutive lengths, [39] for color-critical graphs, [1, 42] for complete bipartite graphs, [51, 28] for planar and outerplanar graphs, [52, 59, 8, 61] for graphs without certain minors, [40] for a spectral Erdős–Stone–Bollobás theorem, [41] for the spectral stability theorem, [11] for a spectral Erdős–Sós theorem, [57] for a spectral Erdős–Pósa theorem, [25] for a spectral Erdős–Rademacher theorem, [26] for a spectral Erdős–Faudree–Rousseau theorem, [9, 58] for friendship graphs, [24, 27] for bowties, [60, 43] for book graphs, [10] for odd wheels and [53, 19] for a characterization of spectral extremal problems for a class of graphs.
Although there has been a wealth of research on the spectral extremal problems in recent years, there are very few spectral results on the problem of counting substructures. In particular, Bollobás and Nikiforov [7] proved that every $n$-vertex graph has at least $(\frac{\lambda}{n}-1+\frac{1}{r})\frac{r(r-1)}{r+1}(\frac{n}{r})^{r+1}$ copies of $K_{r+1}$. This result provides a spectral supersaturation on cliques. Moreover, the spectral supersaturation for a general graph can be found in [25]. Let $F$ be a graph with chromatic number $\chi(F)=r+1$. For any $\varepsilon>0$, there exist $\delta>0$ and $n_{0}\in\mathbb{N}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with $\lambda(G)\geq\left(1-\frac{1}{r}+\varepsilon\right)n$, then $G$ contains at least $\delta n^{f}$ copies of $F$, where $f$ is the number of vertices of $F$.
Recall in Theorem 1.1 that $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles whenever $e(G)>e(T_{n,2})$. Correspondingly, it is natural to consider the spectral version: if $G$ is a graph with $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles. However, this intuition is not true. Let $K_{a,b}^{+}$ be the graph obtained from the complete bipartite graph $K_{a,b}$ by adding an edge to the vertex set of size $a$. For even $n$, we take $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$. It is easy to verify that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles. Fortunately, it was proved by Ning and Zhai [45] that $\lfloor\frac{n}{2}\rfloor-1$ triangles can be guaranteed, which establishes a spectral version of a result of Erdős [18]; see the remark after Theorem 1.1.
Theorem 1.3 (Ning–Zhai, 2023).
If $G$ is a graph on $n$ vertices with
$\lambda(G)\geq\lambda(T_{n,2}),$
then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G$ is the bipartite Turán graph $T_{n,2}$.
Furthermore, Ning and Zhai [45] proposed a general problem on finding spectral counting results for color-critical graphs, including, e.g., the spectral version of the Lovász–Simonovits result in Theorem 1.2. Incidentally, this problem was also selected as one of the “Unsolved Problems” in spectral graph theory [31, Section 9]. To the best of our knowledge, there are few spectral results on counting substructures in a graph with given order, including only triangles [45], cliques [7], triangular edges [26] and bowties [27]. The main contribution of this paper is to establish a spectral version of Theorem 1.2. To some extent, our result (in next section) could be viewed as a partial resolution of the Ning–Zhai problem.
2 Main results
2.1 A spectral Lovász–Simonovits theorem
Before showing a spectral version of Theorem 1.2, the first obstacle one may ask is to establish a spectral condition corresponding to the conventional size-condition. How to determine an increment function $\delta(q)>0$ such that the condition $e(G)\geq\lfloor n^{2}/4\rfloor+q$ can be replaced with the spectral condition $\lambda(G)\geq\lambda(T_{n,2})+\delta(q)$. For instance, is it sufficient to take $\delta(q)=2q/n$? As seen from the phenomenon, where every additional edge guarantees $\lfloor n/2\rfloor$ additional triangles, it is unlikely that such an increment on the spectral radius can create $\lfloor n/2\rfloor$ new triangles. From this point of view, it seems difficult to measure the increment on the spectral radius of a graph. To get around this difficulty, Li, Lu and Peng [25] presented a new way to measure the spectral increment. To start with, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the larger vertex part; see Figure 1. Under this notation, they define the spectral increment as $\delta(q)=\lambda(T_{n,2,q})-\lambda(T_{n,2})$. In particular, they [25, Problem 5.3] proposed the following problem.
Problem 2.1 (Li–Lu–Peng, 2024).
Let $1\leq q<{n}/{2}$ and $G$ be an $n$-vertex graph with
$\lambda(G)\geq\lambda(T_{n,2,q}).$
Then $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G=T_{n,2,q}$.
Problem 2.1 could be regarded as a spectral version of the Lovász–Simonovits theorem. Especially, Li, Lu and Peng [25] confirmed Problem 2.1 in the fundamental case of $q=1$. In the present paper, our goal is to solve this problem for a general $q\in\mathbb{N}$. Let $t(G)$ be the number of triangles in $G$. Moreover, we will show further that $t(G)\geq q\lfloor{n}/{2}\rfloor$ in Problem 2.1 can be guaranteed by a slightly weaker condition. Let $Y_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the vertex part of larger size; see Figure 1. Clearly, we can see that $e(Y_{n,2,q})=e(T_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ and $t(Y_{n,2,q})=t(T_{n,2,q})=q\lfloor n/2\rfloor$. Upon direct computation, it is not difficult to verify that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
Figure 1: The graphs $Y_{n,2,q}$ and $T_{n,2,q}$.
In what follows, we are going to present the spectral version of the Lovász–Simonovits theorem. First of all, we give a result stronger than Problem 2.1.
Theorem 2.2.
If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with
$\lambda(G)\geq\lambda(Y_{n,2,q}),$
then there are at least $q\lfloor{n}/{2}\rfloor$ triangles in $G$.
Remark. The spectral condition in Theorem 2.2 is easier to fulfill than the size condition in Theorem 1.2. Indeed, invoking the fact $\lambda(G)\geq 2e(G)/n$, we can verify that if $e(G)>e(Y_{n,2,q})$, then $\lambda(G)>\lambda(Y_{n,2,q})$. Otherwise, assuming that $\lambda(G)\leq\lambda(Y_{n,2,q})$, it yields that $e(G)\leq\lfloor\frac{n}{2}\lambda(G)\rfloor\leq\lfloor\frac{n}{2}\lambda(Y_{n,2,q})\rfloor=e(Y_{n,2,q})$, where the last equality holds by $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{8q}{n^{2}}$. In conclusion, the spectral condition is weaker than the size condition. Additionally, there are a vast number of graphs $G$ with $\lambda(G)>\lambda(Y_{n,2,q})$ while $e(G)<e(Y_{n,2,q})$.
Note that $\lambda(T_{n,2,q})>\lambda(Y_{n,2,q})$. Applying Theorem 2.2, we know that the desired bound in Problem 2.1 holds immediately. By using the double-eigenvector technique, we shall confirm Problem 2.1 by proving that $T_{n,2,q}$ is the unique graph that attains the maximum spectral radius among all $n$-vertex graphs with at most $q\lfloor{n}/{2}\rfloor$ triangles.
Theorem 2.3.
If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with
$\lambda(G)\geq\lambda(T_{n,2,q}),$
then $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G=T_{n,2,q}$.
When we embed a graph $H$ with $q$ edges into a vertex part of the bipartite Turán graph $T_{n,2}$, we can compute that the spectral radius of the resulting graph will be decreasing in the order of star, complete graph, complete bipartite graph, cycle, path and matching. In addition, we can verify that if we embed two matchings of size $q_{1}$ and $q_{2}$ with $q_{1}+q_{2}=q$ into two vertex parts, respectively, then the resulting graph has larger spectral radius than that of $Y_{n,2,q}$. On the contrary, if we embed two stars of sizes $q_{1}$ and $q_{2}$ into two parts, respectively, then the resulting graph has spectral radius smaller than that of $T_{n,2,q}$.
2.2 A spectral extension on a stability result
In 2021, Xiao and Katona [56] initiated the study on stability of Theorem 1.1. The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices in a set that contains at least one vertex of each triangle in $G$. In particular, $\tau_{3}(G)=1$ if and only if all triangles of $G$ share a common vertex. It was shown in [56] that if $G$ has more than $\lfloor n^{2}/4\rfloor$ edges and $\tau_{3}(G)\geq 2$, then $G$ has at least $n-2$ triangles. Furthermore, they proposed a problem to study the stability of Theorem 1.2. Recently, this problem was solved by Liu and Mubayi [32], as well as Balogh and Clemen [3] independently.
Before stating their result, we need to define two extremal graphs. Firstly, let $G_{1}$ be the graph of order $n$, where $V(G_{1})=A\cup B$ with $|A|=\lceil{n}/{2}\rceil+a$ and $|B|=\lfloor{n}/{2}\rfloor-a$, where $a\in\mathbb{N}$ is chosen later. Pick $2(s-1)$ vertices $x_{1},y_{1},\ldots,x_{s-1},y_{s-1}$ in $A$ and two vertices $u_{1},u_{2}$ in $B$. Adding $s$ edges $\{u_{1},u_{2}\}$ and $\{x_{i},y_{i}\}$ into $K_{|A|,|B|}$, and delete $\alpha$ edges $\{u_{1},x_{1}\},\ldots,\{u_{1},x_{\alpha}\}$, where $\alpha:=s-t-a^{2}-\mathbf{1}_{\{2\nmid n\}}a$. Then $e(G_{1})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{1})=s$. Secondly, let $G_{2}$ be the graph with the same vertex partition. Unlike the previous construction, we now pick $2s$ vertices $x_{1},y_{1},\ldots,x_{s},y_{s}\in A$ and one vertex $u\in B$, add the edges $\{x_{i},y_{i}\}$ into $A$, and delete the edges $\{u,x_{1}\},\ldots,\{u,x_{\alpha}\}$ between $A$ and $B$. Similarly, we have $e(G_{2})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{2})=s$. It is easy to see that $t(G_{1})=(s-1)|B|+|A|-2\alpha$ and $t(G_{2})=s|B|-\alpha$. Now, we choose an appropriate $a\in\mathbb{N}$ to minimize the number of triangles in the above constructions.
Theorem 2.4 (Balogh–Clemen, 2023).
Let $t,s\in\mathbb{N}$ such that $0<t<s$. Suppose that $G$ is an $n$-vertex graph with $\lfloor n^{2}/4\rfloor+t$ edges. If $\tau_{3}(G)\geq s$ and $n$ is sufficiently large, then $G$ contains at least $\min\{t(G_{1}),t(G_{2})\}$ triangles.
An independent work was also studied by Liu and Mubayi [32, Theorem 1.4]. In other words, there are two ways to embed $s$ disjoint edges into the partition in order to minimize the number of triangles, i.e., embedding $s-1$ or $s$ edges into the larger vertex part, and other edges into the smaller part. Recently, Li, Feng and Peng [27] presented a spectral version of the result of Xiao and Katona by showing that if $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq 2$, then $G$ contains at least $n-3$ triangles. This bound is the best possible and it provides a spectral stability of Ning–Zhai’s result in Theorem 1.3.
Next, we shall consider the general case $\tau_{3}(G)\geq s$ for every $s\in\mathbb{N}$. The third main result in this paper reveals an asymptotically spectral extension on Theorem 2.4.
Theorem 2.5.
If $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, then for every $n\geq 113s^{2}$, there are at least $s{n}/{2}-5s^{2}$ triangles in $G$.
Remark. Theorem 2.5 gives a stability of Theorem 2.3. By applying the fact $\lambda(G)\geq 2e(G)/n$, we can see that $e(G)>e(T_{n,2})$ implies $\lambda(G)>\lambda(T_{n,2})$. As a by-product of Theorem 2.5, we obtain that if $e(G)>e(T_{n,2})$ and $\tau_{3}(G)\geq s$, then $G$ has $s{n}/{2}-O(1)$ triangles. This result could be viewed as an asymptotic version of Theorem 2.4, since $t(G_{1})\approx t(G_{2})\approx s{n}/{2}-O(1)$. With additional efforts, one can improve the constant in Theorem 2.5, but it seems difficult to determine the exact bound or the spectral extremal graphs corresponding to the edge-extremal graphs $G_{1}$ and $G_{2}$ in the general case.
Our approach. The main idea in our proofs of Theorems 2.2, 2.3 and 2.5 uses a very simple yet surprisingly powerful technique known as the supersaturation-stability; see Theorem 3.3. Informally, every graph with few triangles can be made bipartite by deleting a quite small number of edges. The word “quite small” means a small constant number of edges, say $6q$ in Lemma 4.1, which is independent of the order $n$. This is slightly different from the classical stability method, which requires us to delete roughly $\varepsilon n^{2}$ edges with a small number $\varepsilon>0$. Although this might sound somewhat vague, we will make it more precise in Section 4. Unlike the traditional stability method, it is sufficient to utilize the condition $n\geq 300q^{2}$ only, instead of requiring that $q$ is fixed and $n$ is sufficiently large (or a tower-type function on $q$), since our proof avoids the use of the celebrated triangle removal lemma***The lemma says that for each $\varepsilon>0$, there is $\delta>0$ such that every graph on $n$ vertices with at most $\delta n^{3}$ triangles can be made triangle-free by removing at most $\varepsilon n^{2}$ edges. That is, a graph with a subcubic number of triangles can be made triangle-free by removing a subquadratic number of edges; see [13].. This allows us to put $q$ in a wide range that can be growth with $n$. Hence, we substantially improve the lower bound on the order $n$ of a graph with respect to $q$. We believe that the supersaturation-stability method used in this paper may be interesting in its own right, and perhaps could have other applications elsewhere, just as the classical stability approach has now seen widespread use. Last but not least, apart from using the supersaturation-stability, we shall make use of the spectral technique due to Cioabă, Feng, Tait and Zhang [9] as well as the double-eigenvector technique applied by Fang, Tait and Zhai [19].
Organization. In Section 3, we shall introduce some preliminaries, including the spectral triangle counting lemma and the supersaturation-stability; see Lemma 3.2 and Theorem 3.3, respectively. Intuitively it is clear that if a graph $G$ has large spectral radius and it is far from being bipartite, then it is unlikely that all the triangles of $G$ are located on few edges of $G$. To warm up, we shall give a “strong” stability result for graphs with few triangles and large spectral radius; see Theorem 3.4. In Section 4, we will outline the sketch of our method and then we present the detailed proof of Theorem 2.2. In Section 5, we shall use the double-eigenvector technique to give the proof of Theorem 2.3. The proof of Theorem 2.5 will be provided in Section 6. At the end of this paper, we will propose some related spectral extremal graph problems for interested readers.
Notation. We usually write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where each $e_{i}$ is a two-element subset of $V$. Here, we admit $|V|=n$ and $|E|=m$. If $S\subseteq V$ is a subset of the vertex set, then $G[S]$ denotes the subgraph of $G$ induced by $S$, i.e., the graph on $S$ whose edges are those edges of $G$ with both endpoints in $S$. For abbreviation, we denote $e(S)=e(G[S])$. We will write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and similarly, we write $e(S,T)$ for the number of edges of $G[S,T]$. For every $s,t\in\mathbb{N}$, we write $K_{s,t}$ for the complete bipartite graph with parts of sizes $s$ and $t$. Let $N(v)$ be the set of vertices adjacent to a vertex $v$ and let $d(v)=|N(v)|$. Moreover, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$ for simplicity. We will write $\lambda(G)$ for the spectral radius of $G$ and $t(G)$ for the number of triangles in $G$. Sometimes, we write $\lambda$ and $t$ for short. The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices that hit all triangles of $G$. For a vertex set $S\subseteq V(G)$, we write $G\setminus S$ for the subgraph of $G$ by deleting all vertices of $S$ and their incident edges. Let $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})^{\top}$ be a Perron vector of $G$, i.e., a non-negative eigenvector corresponding to $\lambda(G)$. By scaling, we may assume further that $\max\{x_{v}:v\in V(G)\}=1$. For a subset $S\subseteq V(G)$, we denote $x_{S}=\sum_{v\in S}x_{v}$ for notational convenience. In particular, we have $x_{V}=\sum_{v\in V(G)}x_{v}$.
3 Preliminaries
Note that in the case $q=1$, we have $Y_{n,2,1}=T_{n,2,1}$.
Lemma 3.1 (See [25]).
(a) If $n$ is even, then $\lambda(Y_{n,2,1})$ is the largest root of
$f(x)=x^{3}-x^{2}-(n^{2}x)/4+{n^{2}}/{4}-n.$
(b) If $n$ is odd, then $\lambda(Y_{n,2,1})$ is the largest root of
$g(x)=x^{3}-x^{2}+{(1-n^{2})x}/{4}+{n^{2}}/{4}-n+{3}/{4}.$
Consequently, we have
$\lambda^{2}(Y_{n,2,1})>\left\lfloor{n^{2}}/{4}\right\rfloor+2.$
In 2007, Bollobás and Nikiforov [7] established some powerful results on counting the number of triangles (cliques) of a graph in terms of the spectral radius.
Lemma 3.2 (See [7, 9, 45]).
Let $G$ be a graph with $m$ edges. Then
$t(G)\geq\frac{\lambda\bigl{(}\lambda^{2}-m\bigr{)}}{3}.$
The equality holds if and only if $G$ is a complete bipartite graph.
The inequality can be written as the following versions:
$\lambda^{3}\leq 3t+m\lambda\quad\Leftrightarrow\quad{t\geq\frac{1}{3}\lambda(\lambda^{2}-m)}\quad\Leftrightarrow\quad m\geq\lambda^{2}-\frac{3t}{\lambda}.$
This inequality was initially obtained by Bollobás and Nikiforov [7, Theorem 1], and it was independently discovered by Cioabă, Feng, Tait and Zhang [9]. The case of equality was recently characterized by Ning and Zhai [45].
Moon–Moser’s inequality states that if $G$ has $n$ vertices and $m$ edges, then
$t(G)\geq\frac{4m}{3n}\left(m-\frac{n^{2}}{4}\right).$
We refer to [4, p. 297] and [33, p. 443] for detailed proofs. This result gives the supersaturation on triangles for graphs with more than $n^{2}/4$ edges. Next, we present a generalization for a graph with less than $n^{2}/4$ edges whenever this graph is far from being bipartite. More precisely, for $\varepsilon>0$, we say that a graph $G$ is $\varepsilon$-far from being bipartite if $G^{\prime}$ is not bipartite for every subgraph $G^{\prime}\subseteq G$ with $e(G^{\prime})>e(G)-\varepsilon$. In other words, if $G$ is $\varepsilon$-far from being bipartite, then we need to delete at least $\varepsilon$ edges from $G$ to make it bipartite.
Theorem 3.3 (See [2]).
If $G$ is $\varepsilon$-far from being bipartite, then
$t(G)\geq\frac{n}{6}\left(m+\varepsilon-\frac{n^{2}}{4}\right).$
The supersaturation on triangles is usually deduced from Moon–Moser’s inequality. In fact, Theorem 3.3 can also imply the supersaturation on triangles. Note that every bipartite graph of order $n$ has at most $n^{2}/4$ edges. Thus, if $G$ has at least ${n^{2}}/{4}+q$ edges, then it is $q$-far from being bipartite. Theorem 3.3 implies that $G$ contains at least ${qn}/{3}$ triangles. In particular, setting $q=\varepsilon n^{2}$, we obtain that $G$ has at least ${\varepsilon n^{3}}/{3}$ triangles.
For dealing with some extremal graph problems, it is a great advantage that applying Theorem 3.3 can help us to get rid of the use of triangle removal lemma so that we do not require the order of a graph to be sufficiently large. This approach could be considered to be a key innovation in our paper, and it was proven to be effective and feasible in many extremal graph problems. We refer to [26, 27] for more related applications. In what follows, we provide a warm-up practice that has its independent interest. We concentrate on explaining the main ideas which we believe have wide applicability.
Theorem 3.4.
If $G$ is a graph on $n$ vertices with $\lambda(G)\geq n/2$ and $t(G)\leq kn/2$, then
$e(G)\geq\frac{n^{2}}{4}-3k,$
and there exists a vertex partition $V(G)=S\cup T$ such that
$e(S,T)\geq\frac{n^{2}}{4}-9k$
and
$\frac{n}{2}-3\sqrt{k}\leq|S|,|T|\leq\frac{n}{2}+3\sqrt{k}.$
Moreover, we have
$\frac{n}{2}-12k\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+9k.$
Theorem 3.4 gives an effective way to obtain a bipartition of a graph with large spectral radius and few triangles. To some extend, this theorem indicates how the supersaturation-stability method can be used to deal with a variety of spectral extremal problems. The proof of Theorem 3.4 is a well combination of Lemma 3.2 and Theorem 3.3.
Proof.
First of all, if $G$ is a complete bipartite graph, then $e(G)\leq\lfloor n^{2}/4\rfloor$. Moreover, we have $e(G)=\lambda^{2}\geq n^{2}/4$. It follows that $n$ is even and $G=K_{\frac{n}{2},\frac{n}{2}}$. In this case, the desired result holds immediately. Now, we consider the case that $G$ is not a complete bipartite graph, then the inequality in Lemma 3.2 holds strictly. So we get $e(G)>\lambda^{2}-(3t)/\lambda\geq\lambda^{2}-(6t)/n\geq n^{2}/4-3k$. Theorem 3.3 implies that $G$ is not $6k$-far from being bipartite. Otherwise, if $G$ is $6k$-far from being bipartite, then $t(G)\geq{n}/{6}\cdot(e(G)+6k-n^{2}/4)>kn/2$, which contradicts with the assumption. Hence $G$ is not $6k$-far from being bipartite. Then there exists a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6k$. Moreover, we can get $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4-9k$. Consequently, we obtain that $n/2-3\sqrt{k}<|S|,|T|<n/2+3\sqrt{k}$. Note that $e(K_{|S|,|T|})-e(S,T)<|S||T|-(n^{2}/4-9k)\leq 9k$. In other words, there are less than $9k$ missing edges between $S$ and $T$. Consequently, we get $\delta(G)>(n/2-3\sqrt{k})-9k\geq n/2-12k$. Note that $e(S)+e(T)<6k$. Then it follows that $\Delta(G)<(n/2+3\sqrt{k})+6k\leq n/2+9k$, as desired. ∎
4 Proof of Theorem 2.2
Note that the case $q=1$ in Problem 2.1 was proved by Li, Lu and Peng [25]. Next, we shall consider the case $q\geq 2$ and $n\geq 300q^{2}$ although our proof works for $q=1$. To avoid unnecessary computations, we did not make much efforts to optimize the bound on the order of a graph. Throughout this section, let $G$ be a graph on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$. Suppose on the contrary that $G$ contains less than $q\lfloor{n}/{2}\rfloor$ triangles. In the sequel, we will deduce a contradiction. The essential observation in our proof attributes to the supersaturation-stability as described in Theorem 3.3.
To start the proof of Theorem 2.2, we outline the main steps as follows.
  • Firstly, we establish a bipartition of vertices of $G$. Using Lemma 3.2, we get that $G$ has more than $n^{2}/4-3q$ edges. The supersaturation-stability in Theorem 3.3 implies that one can delete $6q$ edges from $G$ to make it bipartite. So we get a partition $V(G)=S\cup T$ with $e(S)+e(T)<6q$ and $|S|,|T|\approx n/2\pm 3\sqrt{q}$.
  • Secondly, we shall show further that $e(S)+e(T)\leq q$. Moreover, the minimum degree of $G$ is at least $n/2-4q$; see Lemmas 4.2 and 4.3. Let $\mathbf{x}\in\mathbb{R}^{n}$ be the Perron vector of $G$ satisfying $\max\{x_{v}:v\in V(G)\}=1$. Then we will prove in Lemma 4.4 that ${x}_{u}\geq 1-{30q}/{n}$ for every $u\in V(G)$.
  • Finally, we prove in Lemma 4.6 that the bipartition of $G$ is balanced, i.e., $\bigl{|}|S|-|T|\bigr{|}\leq 1$. Using the above lemmas, we shall prove that $G$ has at least $q\lfloor n/2\rfloor$ triangles, or $G$ satisfies $\lambda(G)<\lambda(Y_{n,2,q})$. This leads to a contridiction.
In what follows, we divide the proof into a sequence of lemmas.
Lemma 4.1.
There is a partition $V(G)=S\cup T$ such that
$e(S)+e(T)<6q$
and
$e(S,T)>\frac{n^{2}}{4}-9q.$
Consequently, we get
$\frac{n}{2}-3\sqrt{q}<|S|,|T|<\frac{n}{2}+3\sqrt{q}.$
Proof.
First of all, we know from Lemma 3.1 that $\lambda^{2}(Y_{n,2,1})>\lfloor n^{2}/4\rfloor+2$. Then $\lambda^{2}(G)\geq\lambda^{2}(Y_{n,2,q})>\lfloor n^{2}/4\rfloor+2$. Since $t(G)<q\lfloor{n}/{2}\rfloor$, by Lemma 3.2, we have
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n}>\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q.$ (1)
We claim that $G$ is not $6q$-far from being bipartite. Otherwise, if $G$ is $6q$-far from being bipartite, then using Theorem 3.3, we get
$t(G)\geq\frac{n}{6}\left(e(G)+6q-\frac{n^{2}}{4}\right)>q\left\lfloor\frac{n}{2}\right\rfloor,$
which contradicts with the assumption. So $G$ is not $6q$-far from being bipartite. That is to say, we can delete less than $6q$ edges from $G$ to make it bipartite, i.e., there exists a vertex partition $V(G)=S\cup T$ such that
$e(S)+e(T)<6q.$
From the above discussion, we get
$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-9q.$
Without loss of generality, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq{n}/{2}-3\sqrt{q}$. Then $|T|=n-|S|\geq{n}/{2}+3\sqrt{q}$. It yields that
$e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-3\sqrt{q}\right)\left(\frac{n}{2}+3\sqrt{q}\right)=\frac{n^{2}}{4}-9q,$
a contradiction. Thus, we must have ${n}/{2}-3\sqrt{q}<|S|,|T|<{n}/{2}+3\sqrt{q}$. ∎
Lemma 4.2.
We have $e(S)+e(T)\leq q$.
Proof.
Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$. Suppose on the contrary that $e(S)+e(T)\geq q+1$. We write $K_{|S|,|T|}$ for the complete bipartite graph on the vertex parts $S$ and $T$. Note that each edge of $G[S]$ or $G[T]$ can yield at least $\min\{|S|,|T|\}$ triangles in $K_{|S|,|T|}$. Recall in Lemma 4.1 that $e(S)+e(T)<6q$. So each missing edge between $S$ and $T$ destroys at most $6q$ triangles. By Lemma 4.1 again, we have $\min\{|S|,|T|\}>{n}/{2}-3\sqrt{q}$, and $G[S,T]$ misses at most $9q$ edges from the graph $K_{|S|,|T|}$. Hence, we conclude
$t(G)\geq(q+1)\left(\frac{n}{2}-3\sqrt{q}\right)-9q\cdot 6q>(q+1)\frac{n}{2}-58q^{2}\geq q\left\lfloor\frac{n}{2}\right\rfloor,$
where the last inequality holds for every $n\geq 116q^{2}$. This contradicts with the assumption $t(G)<q\lfloor n/2\rfloor$. Thus, we have $e(S)+e(T)\leq q$. ∎
Lemma 4.3.
We have
$e(S,T)>\frac{n^{2}}{4}+2-4q$
and
$\frac{n}{2}-2\sqrt{q}<|S|,|T|<\frac{n}{2}+2\sqrt{q}.$
Furthermore, it follows that
$\frac{n}{2}-4q\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+q+2\sqrt{q}.$
Proof.
By Lemma 4.2, we obtain $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4+2-4q$. By the AM-GM inequality, we can get ${n}/{2}-2\sqrt{q}\leq|S|,|T|\leq{n}/{2}+2\sqrt{q}$.
By Rayleigh’s formula, it is well-known that $\delta(G)\leq\lambda(G)\leq\Delta(G)$. Since $e(S)+e(T)\leq q$, each vertex of $S$ has degree at most $q$ in $G[S]$, and this also holds for the vertices of $T$. Then it follows that $\Delta(G)\leq\max\{|S|,|T|\}+q\leq{n}/{2}+2\sqrt{q}+q$. Next, we show that $\delta(G)\geq{n}/{2}-4q$. Otherwise, if there exists a vertex $v\in V(G)$ with the minimum degree and $d(v)\leq{n}/{2}-4q-{1}/{2}$, then
$e(G\setminus\{v\})=e(G)-d(v)$$\geq$ $\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q-\left(\frac{n}{2}-4q-\frac{1}{2}\right)$
$\geq$ $\frac{(n-1)^{2}}{4}+q+2.$
By Theorem 1.2, we can find $(q+2)\lfloor\frac{n-1}{2}\rfloor$ triangles in $G\setminus\{v\}$. This is a contradiction with the assumption $t(G)<q\lfloor{n}/{2}\rfloor$. Thus, we must have $\delta(G)\geq{n}/{2}-4q$. ∎
Recall that $\mathbf{x}\in\mathbb{R}^{n}$ is the Perron vector of $G$ with $\max\{x_{v}:v\in V(G)\}=1$.
Lemma 4.4.
For every $u\in V(G)$, we have
${x}_{u}>1-\frac{30q}{n}.$
Proof.
Let $z\in V(G)$ be a vertex such that ${x}_{z}$ attains the maximum eigenvector entry of $\mathbf{x}$ and ${x}_{z}=1$. Without loss of generality, we may assume that $z\in S$. Our proof of this lemma falls naturally into two steps.
Step 1. For every $u\in S$, we have ${x}_{u}>1-\frac{15q}{n}$.
Since $e(S)\leq q$ by Lemma 4.2, we get $d_{S}(u)\leq q$. By Lemma 4.3, we have $d_{T}(u)=d(u)-d_{S}(u)\geq({n}/{2}-4q)-q={n}/{2}-5q$. By Lemma 4.3 again, we get $|T|\leq{n}/{2}+2\sqrt{q}$. Then
$|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|$ $=$ $-|N_{T}(u)|+|N_{T}(u)\cup N_{T}(z)|$
$\leq$ $-\left(\frac{n}{2}-5q\right)+\left(\frac{n}{2}+2\sqrt{q}\right)$
$=$ $5q+2\sqrt{q}.$
Hence, we have
$\lambda{x}_{u}-\lambda{x}_{z}=\sum_{v\sim u}{x}_{v}-\sum_{v\sim z}{x}_{v}\geq-\sum_{v\sim z,v\in T,v\not\sim u}{x}_{v}-\sum_{v\sim z,v\in S}{x}_{v}.$
Applying the fact ${x}_{v}\leq 1$, it follows that
$\lambda{x}_{u}-\lambda{x}_{z}\geq-\Bigl{(}|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|\Bigr{)}-|N_{S}(z)|\geq-6q-2\sqrt{q}.$
Recall that ${x}_{z}=1$ and $\lambda>{n}/{2}$. Then for each $u\in S$, we have
${x}_{u}\geq 1-\frac{6q+2\sqrt{q}}{\lambda}>1-\frac{12q+4\sqrt{q}}{n}>1-\frac{15q}{n}.$
Step 2. For every $u\in T$, we have ${x}_{u}>1-\frac{30q}{n}$.
By the previous step, we get
$\lambda{x}_{u}=\sum_{v\in N(u)}{x}_{v}\geq\sum_{v\in N_{S}(u)}{x}_{v}\geq\left(1-\frac{12q+4\sqrt{q}}{n}\right)d_{S}(u).$
By Lemma 4.3, we can see that $d(u)\geq\delta(G)\geq{n}/{2}-4q$, which together with $d_{T}(u)\leq q$ gives $d_{S}(u)=d(u)-d_{T}(u)\geq{n}/{2}-5q$. Hence, for $q\geq 2$, we have
${x}_{u}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})d_{S}(u)}{\lambda}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})(\frac{n}{2}-5q)}{\frac{n}{2}+q+2\sqrt{q}}>1-\frac{30q}{n}.$
Combining the above two claims, the desired result follows immediately. ∎
Using Lemma 4.4, namely, ${x}_{u}=1-O({1}/{n})$, we will show that the bipartition $V(G)=S\cup T$ is balanced (see the forthcoming Lemma 4.6). To start with, we fix some notation for convenience. Let $K_{|S|,|T|}$ be the complete bipartite graph with partite sets $S$ and $T$. Let $G_{1}=G[S]\cup G[T]$ be the graph whose edges consist of the class-edges, and let $G_{2}$ be the graph on $V(G)$ with the missing edges between $S$ and $T$, that is, $E(G_{2})=E(K_{|S|,|T|})\setminus E(G)$. Note that $e(G)=e(G_{1})+e(K_{|S|,|T|})-e(G_{2})$.
Lemma 4.5.
We have
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}\leq\frac{120q^{2}}{n(n-60q)}.$
Proof.
By Lemma 4.2, we get $e(G_{1})\leq q$. By Lemma 4.4, we have
$\mathbf{x}^{\top}\mathbf{x}>n\left(1-\frac{30q}{n}\right)^{2}>n\left(1-\frac{60q}{n}\right)=n-60q.$
Using Rayleigh’s formula, it follows that
$\lambda(G)$ $=$ $\frac{\mathbf{x}^{\top}\bigl{(}A(K_{|S|,|T|})+A(G_{1})-A(G_{2})\bigr{)}\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
$=$ $\frac{\mathbf{x}^{\top}A(K_{|S|,|T|})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}+\frac{\mathbf{x}^{\top}A(G_{1})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
$<$ $\sqrt{|S||T|}+\frac{2q}{n-60q}.$
On the other hand, it is familiar that
$\lambda(G)\geq\lambda(Y_{n,2,q})\geq\frac{2e(Y_{n,2,q})}{n}\geq\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor+\frac{2q}{n}.$
Then we have
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<2q\left(\frac{1}{n-60q}-\frac{1}{n}\right)=\frac{120q^{2}}{n(n-60q)}.$
This completes the proof. ∎
Lemma 4.6.
We have $\bigl{|}|S|-|T|\bigr{|}\leq 1$.
Proof.
We assume on the contrary that $|T|\geq|S|+2$. Next, considering the following two cases, and we will derive a contradiction in each case.
Case 1. $n$ is even. Since $|S|+|T|=n$, we get
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}$$\geq$ $\frac{n}{2}-\sqrt{\left(\frac{n}{2}-1\right)\left(\frac{n}{2}+1\right)}$
$=$ $\frac{n}{2}-\sqrt{\frac{n^{2}}{4}-1}=\frac{1}{\frac{n}{2}+\sqrt{\frac{n^{2}}{4}-1}}>\frac{1}{n}.$
Case 2. $n$ is odd. Since $|S|+|T|=n$, we have
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}$$\geq$ $\frac{n^{2}-1}{2n}-\sqrt{\left(\frac{n-3}{2}\right)\left(\frac{n+3}{2}\right)}$
$=$ $\frac{1}{2}\left({n-\frac{1}{n}}-{\sqrt{n^{2}-9}}\right)=\frac{7+\frac{1}{n^{2}}}{2(n-\frac{1}{n}+\sqrt{n^{2}-9})}>\frac{1}{n}.$
By Lemma 4.5, we get
$\frac{1}{n}<\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{120q^{2}}{n(n-60q)}.$
This leads to a contradiction for $q\geq 2$ and $n\geq 150q^{2}$. ∎
Recall that $G$ is an $n$-vertex graph with less than $q\lfloor{n}/{2}\rfloor$ triangles. In what follows, we will deduce a contradiction in two cases. On the one hand, if $G$ has at least $\lfloor n^{2}/4\rfloor+q$ edges, then Theorem 1.2 implies that $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles, which is a contradiction with the assumption. On the other hand, if $G$ has at most $\lfloor n^{2}/4\rfloor+q-1$ edges, then we will prove that $\lambda(G)<\lambda(Y_{n,2,q})$, which leads to a contradiction as well. This will completes the proof of Theorem 2.2. We write the second case as the following lemma.
Lemma 4.7.
If $e(G)\leq\lfloor n^{2}/4\rfloor+q-1$, then $\lambda(G)<\lambda(Y_{n,2,q})$.
Proof.
By Lemma 4.6, we know that $\bigl{|}|S|-|T|\bigr{|}\leq 1$. Recall that $Y_{n,2,q}$ is the graph on the vertex sets $S$ and $T$ with all edges between $S$ and $T$, together with $q$ pairwise disjoint edges embedded into the vertex part of size $\lceil{n}/{2}\rceil$. Clearly, we have $e(Y_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ edges. We denote
$E_{+}:=E(Y_{n,2,q})\setminus E(G)$
and
$E_{-}:=E(G)\setminus E(Y_{n,2,q}).$
Note that $(E(G)\cup E_{+})\setminus E_{-}=E(Y_{n,2,q})$. It follows that
$e(G)+|E_{+}|-|E_{-}|=e(Y_{n,2,q}).$
Using the assumption $e(G)\leq\lfloor n^{2}/4\rfloor+q-1$, we get
$|E_{+}|\geq|E_{-}|+1.$
Furthermore, we know from Lemma 4.2 that $|E_{-}|\leq e(S)+e(T)\leq q$. By Lemma 4.2 again, we have that $|E_{+}|\leq\lfloor{n^{2}}/{4}\rfloor-e(S,T)+q\leq 5q$. By Lemma 4.4, we know that ${x}_{u}\geq 1-{(30q)}/{n}$ for every $u\in V(G)$. Therefore, we obtain
$\lambda(Y_{n,2,q})$$\geq$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(Y_{n,2,q})}{x}_{i}{x}_{j}$
$=$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(G)}{x}_{i}{x}_{j}+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{+}}{x}_{i}{x}_{j}-\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{-}}{x}_{i}{x}_{j}$
[…]
$=$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(|E_{+}|-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|-|E_{-}|\right)$
$\geq$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(1-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|\right)$
$>$ $\lambda(G),$
where the last inequality holds for $n\geq 300q^{2}$ since $|E_{+}|\leq 5q$. Therefore, we have $\lambda(Y_{n,2,q})>\lambda(G)$, as desired. ∎
From the above proof of Lemma 4.7, we can obtain the following proposition.
Proposition 4.8.
If $n\geq 300q^{2}$ and $G$ is a graph obtained from $T_{n,2}$ by adding $q-1$ edges into the color classes arbitrarily, then $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)<\lambda(Y_{n,2,q})$.
5 Proof of Theorem 2.3
In this section, we give the proof of Theorem 2.3. Let $G$ be an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2,q})$ and $G$ has at most $q\lfloor{n}/{2}\rfloor$ triangles. Among such graphs, we choose $G$ as the graph with the maximum spectral radius. Our aim is to prove that $G$ is the expected extremal graph $T_{n,2,q}$. Note that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$. Then Theorem 2.2 implies that $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles. Similarly, the previous Lemma 4.1 – Lemma 4.6 still hold in this setting. By Lemma 4.7, it follows that $e(G)=\lfloor n^{2}/4\rfloor+q$. By Lemma 4.2, we have $e(S)+e(T)=q$ and $e(S,T)=\lfloor n^{2}/4\rfloor$. Consequently, we know that $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles. Moreover, $G$ is obtained from the bipartite Turán graph $T_{n,2}$ by embedding exactly $q$ edges into the color classes, and there is no triangle among these $q$ edges. Combining these facts, it remains to show that the $q$ edges induce a star in the vertex part of size $\lceil{n}/{2}\rceil$.
Observe that $G$ does not contain $q+1$ pairwise edge-disjoint triangles. In particular, if $n$ is even and $q\neq 3$, then $|S|=|T|={n}/{2}$ and setting $k:=q+1$ in [29, Theorem 7], we know that $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$. However, the proof [29] follows the line of that of forbidding friendship graph in [9], where the triangle removal lemma is used so that it requires that $q$ is fixed and $n$ is sufficiently large. Actually, the result seems ambiguous whenever $q$ is large (say, when $q\geq\log n$). The authors [26] recently provided an alternative new method that avoids the use of triangle removal lemma.
From the above discussion, one of the main difficulties of proving Theorem 2.3 lies in giving a unique characterization of the spectral extremal graph. In the sequel, we shall give a complete proof for every $n\geq 300q^{2}$, instead of for $n$ being large enough. The key ingredient relies on the double-eigenvector technique, which was initially used by Rowlinson [50]. As far as we know, this technique turns out to be effective for treating spectral extremal graph problems; see [58, 29, 19] and references therein. The idea in our proof has its root in [29, 19]. To begin with, we need to introduce two lemmas. The first lemma provides an operation of a connected graph that increases the adjacency spectral radius strictly.
Lemma 5.1 (See [55]).
Assume that $G$ is a connected graph with $u,v\in V$ and $w_{1},\ldots,w_{s}\in N(v)\setminus N(u)$. Let $\mathbf{x}=(x_{1},\ldots,x_{n})^{\top}$ be the Perron vector with $x_{v}$ corresponding to the vertex $v$. Let $G^{\prime}=G-\{vw_{i}:1\leq i\leq s\}+\{uw_{i}:1\leq i\leq s\}$. If ${x}_{u}\geq{x}_{v}$, then $\lambda(G^{\prime})>\lambda(G)$.
From the above lemma, we can get the following corollary.
Corollary 5.2.
For any $q\geq 2$, we have $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
The second lemma gives an upper bound on the sum of squares of degrees in a graph with given size. Here, we shall provide an alternative short proof.
Lemma 5.3 (See [14, 46]).
If $G$ is a graph with $m$ edges, then
$\sum_{v\in V(G)}d^{2}(v)\leq m^{2}+m.$
Proof.
For each $\{u,v\}\in E(G)$, we have $d(u)+d(v)\leq m+1$. Then
$\sum_{v\in V(G)}d^{2}(v)=\sum_{\{u,v\}\in E(G)}(d(u)+d(v))\leq m(m+1).\qed$
Recall that we have shown that $G$ is obtained from $T_{n,2}$ by embedding $q$ edges into its partite sets. Without loss of generality, we may assume that $|S|\leq|T|$. Then $|S|=\lfloor{n}/{2}\rfloor$ and $|T|=\lceil{n}/{2}\rceil$. For notational convenience, we denote $x_{S}:=\sum_{u\in S}x_{u}$ and $x_{T}:=\sum_{u\in T}x_{u}$. Let $H_{1}$ and $H_{2}$ be the subgraphs that are embedded into the vertex parts $S$ and $T$, respectively. To maximize $\lambda(G)$, we need to prove that $H_{1}=\varnothing$ and $H_{2}=K_{1,q}$.
Lemma 5.4.
For each $i=1,2$, if $H_{i}\neq\varnothing$, then $H_{i}$ is a star or a $4$-cycle.
Proof.
Without loss of generality, we shall show that if $H_{1}\neq\varnothing$, then $H_{1}$ is a star. We choose $v^{*}\in S$ as a vertex such that ${x}_{v^{*}}=\max\{{x}_{v}:v\in S\}$. Clearly, the vertex $v^{*}$ can not be an isolated vertex of $G[S]$. First of all, we consider the case $1\leq e(H_{1})\leq 3$. By enumerating, we have $H_{1}\in\{K_{2},2K_{2},K_{1,2},3K_{2},P_{3}\cup K_{2},P_{4},K_{1,3}\}$. We claim that $v^{*}$ is a dominating vertex of $H_{1}$. Otherwise, if there is a vertex $u\in V(H_{1})\setminus N(v^{*})$ and an edge $\{u,v\}\in E(H_{1})$, then let $G^{\prime}$ be the graph obtained from $G$ by deleting the edge $\{u,v\}$, and adding the edge $\{u,v^{*}\}$. Observe that $G^{\prime}$ has at most $q\lfloor{n}/{2}\rfloor$ triangles. By Lemma 5.1, we get $\lambda(G^{\prime})>\lambda(G)$, a contradiction. Thus $v^{*}$ is a dominating vertex. So $H_{1}$ is a star.
Now, we consider the case $e(H_{1})\geq 4$. Similar to the above argument, we can move the edges incident to $v^{*}$. Hence we can obtain that $d_{H_{1}}(v^{*})\geq 3$, unless $H_{1}\cong C_{4}$. For the former case, we need to show that $H_{1}$ is a star. We denote $a=d_{H_{1}}(v^{*})\geq 3$ and $H_{1}^{\prime}=H_{1}\setminus\{v^{*}\}$. If $e(H_{1}^{\prime})=0$, then $H_{1}$ is a star, so we are done. If $b:=e(H_{1}^{\prime})\geq 1$, then we define $G^{\prime}$ as the graph obtained from $G$ by deleting all $b$ edges in $H_{1}^{\prime}$, and then adding $b$ new edges from $v^{*}$ to $w_{1},\ldots,w_{b}\in S\setminus V(H_{1})$. We can see that the edges in $G^{\prime}[S]$ form a star $K_{1,a+b}$, and $G^{\prime}$ contains at most $q\lfloor{n}/{2}\rfloor$ triangles. Thus, we have $\lambda(G^{\prime})\leq\lambda(G)$. We denote $\lambda=\lambda(G)$ and $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$. Let $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})^{\top}$ be the Perron vector of $G^{\prime}$. By the symmetry, we have $y_{v}=y_{w_{1}}$ for every vertex $v\in V(K_{1,a+b})\backslash\{v^{*}\}$. Thus, we have $\lambda^{\prime}y_{v^{*}}=y_{T}+(a+b)y_{w_{1}}$ and $\lambda^{\prime}y_{w_{1}}=y_{T}+y_{v^{*}}$, where $y_{T}=\sum_{u\in T}y_{u}$. Then
$y_{v^{*}}=\frac{\lambda^{\prime}+a+b}{\lambda^{\prime}+1}y_{w_{1}}\geq\frac{\lambda+a+b}{\lambda+1}y_{w_{1}}.$
Using the double-eigenvector technique, we have
$\mathbf{x}^{\top}\mathbf{y}\left(\lambda^{\prime}-\lambda\right)$ $=$ $\mathbf{x}^{\top}\left(A(G^{\prime})-A(G)\right)\mathbf{y}$ (2)
$=$ $\sum_{\{u,v\}\in E(G^{\prime})}(x_{u}y_{v}+x_{v}y_{u})-\sum_{\{u,v\}\in E(G)}(x_{u}y_{v}+x_{v}y_{u})$
$=$ $\sum_{j=1}^{b}\left(x_{v^{*}}y_{w_{j}}+x_{w_{j}}y_{v^{*}}\right)-\sum_{\{u,v\}\in E(H_{1}^{\prime})}\left(x_{u}y_{v}+x_{v}y_{u}\right)$
$\geq$ $b\cdot x_{v^{*}}y_{w_{1}}+b\cdot\frac{\lambda+a+b}{\lambda+1}x_{w_{1}}y_{w_{1}}-\sum_{\{u,v\}\in E(H_{1}^{\prime})}\left(x_{u}+x_{v}\right)y_{w_{1}},$
where we used the fact $x_{w_{1}}=\cdots=x_{w_{b}}$ and $y_{v}=y_{w_{1}}=\cdots=y_{w_{b}}$ for any $v\in V(H_{1}^{\prime})$.
Next, we are going to prove
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0.$
Since $N_{G}(v^{*})=T\cup N_{H_{1}}(v^{*})$, we have
$\lambda x_{v^{*}}\geq x_{T}+ax_{w_{1}}.$ (3)
Moreover, as $N_{G}(w_{1})=T$ and $\lambda x_{w_{1}}=x_{T}$, we have
$\frac{\lambda+a+b}{\lambda+1}\lambda x_{w_{1}}\geq\left(\lambda+a+b-\frac{3}{2}\right)x_{w_{1}}=x_{T}+\left(a+b-\frac{3}{2}\right)x_{w_{1}}.$ (4)
Using the double counting, we get
$\sum_{\{u,v\}\in E(H_{1}^{\prime})}\lambda\left(x_{u}+x_{v}\right)=\sum_{v\in V(H_{1}^{\prime})}d_{H_{1}^{\prime}}(v)\lambda x_{v}.$
For each $v\in V(H_{1}^{\prime})$, we have $\lambda x_{v}\leq x_{T}+(d_{H_{1}^{\prime}}(v)+1)x_{v^{*}}$. By Lemma 5.3, it follows that $\sum_{v\in V(H_{1}^{\prime})}d_{H_{1}^{\prime}}^{2}(v)\leq b^{2}+b$. Consequently, we get
$\sum_{\{u,v\}\in E(H_{1}^{\prime})}\lambda(x_{u}+x_{v})$$\leq$ $\sum_{v\in V(H_{1}^{\prime})}d_{H_{1}^{\prime}}(v)x_{T}+\sum_{v\in V(H_{1}^{\prime})}\left(d_{H_{1}^{\prime}}^{2}(v)+d_{H_{1}^{\prime}}(v)\right)x_{v^{*}}$ (5)
$\leq$ $2bx_{T}+(b^{2}+3b)x_{v^{*}}.$
Combining the above inequalities, we can obtain
[…]
$=$ $b(2a+b-\tfrac{3}{2})x_{w_{1}}y_{w_{1}}-b(b+3)x_{v^{*}}y_{w_{1}}.$
To obtain a contradiction, we would like to show that $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$. It is sufficient to prove that $\frac{x_{v^{*}}}{x_{w_{1}}}<\frac{2a+b-{3}/{2}}{b+3}$. Note that $\lambda x_{v^{*}}=x_{T}+\sum_{u\in N_{H_{1}}(v^{*})}x_{u}\leq x_{T}+ax_{v^{*}}$. Then $x_{v^{*}}\leq\frac{x_{T}}{\lambda-a}$. Observe that $x_{w_{1}}=\frac{x_{T}}{\lambda}$. Therefore, it yields that
$\frac{x_{v^{*}}}{x_{w_{1}}}\leq\frac{\lambda}{\lambda-a}.$
Since $a\geq 3$, we get $(2a+b-\frac{3}{2})-(b+3)=2a-\frac{9}{2}>1$. Recall that $\lambda(G)\geq\lambda(T_{n,2,q})>{n}/{2}$ and $b\leq q-3$. Then for $n\geq 2(q^{2}+q)$, it follows that
$\frac{x_{v^{*}}}{x_{w_{1}}}\leq\frac{\lambda}{\lambda-a}<1+\frac{1}{b+3}<\frac{2a+b-{3}/{2}}{b+3},$
which leads to $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$ and $\lambda^{\prime}>\lambda$, a contradiction. Thus, $H_{1}$ is a star. ∎
In Lemma 5.4, we expect to prove that $H_{i}$ is a star, instead of a $4$-cycle. But the above argument seems invalid for $H_{1}\cong C_{4}$, so we need to make an easy comparison between these two cases. If $n$ is odd, then $|S|={(n-1)}/{2}$ and $|T|={(n+1)}/{2}$. Since $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles, we know that the $q$ edges must be embedded into the larger part $T$. By Lemma 5.4, we must embed a star $K_{1,q}$ or a cycle $C_{4}$ (in the case of $q=4$) into $T$. By direct computation, one can check that embedding $K_{1,4}$ to $T_{n,2}$ attains the maximum spectral radius. Indeed, we can compute that $\lambda(T_{n,2,4})$ is the largest root of
$t(x):=9-10n+n^{2}+4x-4nx-(15x^{2})/4-(n^{2}x^{2})/4+x^{4}.$
Let $G$ be the graph obtained from $T_{n,2}$ by embedding a $C_{4}$ into $T$. Then $\lambda(G)$ is the largest root of
$g(x):=7/2-4n+n^{2}/2+x/4-(n^{2}x)/4-2x^{2}+x^{3}.$
It is not hard to check that $\lambda(G)<\lambda(T_{n,2,4})$. This completes the proof for odd $n$.
If $n$ is even, then $|S|=|T|={n}/{2}$. Similarly, the embedded graphs $H_{1}$ and $H_{2}$ are stars or $4$-cycles by Lemma 5.4. The remaining task is to show that embedding only one star into $S$ or $T$ will lead to a graph with the maximum spectral radius.
Lemma 5.5.
Let $n$ be even and $G^{\prime}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the part $S$. Let $\lambda^{\prime}=\lambda(G^{\prime})$ and $\mathbf{y}$ be the Perron vector of $G^{\prime}$. Then
$\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|S|+\frac{2q}{\lambda^{\prime}-q}}\leq y_{T}\leq\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|S|+\frac{2q}{\lambda^{\prime}}}.$
Proof.
By definition, we have $y_{V}=y_{S}+y_{T}$. It is sufficient to prove that
$|S|y_{T}+2q\cdot\frac{y_{T}}{\lambda^{\prime}}\leq\lambda^{\prime}y_{S}\leq|S|y_{T}+2q\cdot\frac{y_{T}}{\lambda^{\prime}-q}.$
For each $v\in S$, we have $\lambda^{\prime}y_{v}=y_{T}+\sum_{u\in N_{S}(v)}y_{u}$. Then
$\lambda^{\prime}y_{S}=\sum_{v\in S}\lambda^{\prime}y_{v}=|S|y_{T}+\sum_{v\in S}\sum_{u\in N_{S}(v)}y_{u}=|S|y_{T}+\sum_{u\in S}d_{S}(u)y_{u}.$
Note that $\lambda^{\prime}y_{u}\geq y_{T}$ for each $u\in S$. Then
$\min_{u\in S}\{y_{u}\}\geq\frac{y_{T}}{\lambda^{\prime}}.$
We denote $y_{v^{*}}=\max_{u\in S}\{y_{u}\}$. Similarly, we have $\lambda^{\prime}y_{v^{*}}=y_{T}+\sum_{w\in N_{S}(v^{*})}y_{w}\leq y_{T}+d_{S}(v^{*})y_{v^{*}}=y_{T}+qy_{v^{*}}$, which yields $y_{v^{*}}\leq\frac{y_{T}}{\lambda^{\prime}-q}$. Thus, it follows that
$|S|y_{T}+2e(S)\frac{y_{T}}{\lambda^{\prime}}\leq\lambda^{\prime}y_{S}\leq|S|y_{T}+2e(S)\frac{y_{T}}{\lambda^{\prime}-q}.$
The desired inequality holds immediately by $e(S)=e(K_{1,q})=q$. ∎
Lemma 5.6.
There exists exactly one $i\in\{1,2\}$ such that $H_{i}\neq\varnothing$ and $H_{i}\cong K_{1,q}$.
Proof.
Suppose on the contrary that both $H_{1}$ and $H_{2}$ are non-empty. By Lemma 5.4, we know that $H_{i}$ is a star or a $4$-cycle. Thus, we have the following three cases:
  • (a)
    $H_{1}\cong K_{1,q_{1}}$ and $H_{2}\cong K_{1,q_{2}}$, where $q_{1}+q_{2}=q$.
  • (b)
    $H_{1}\cong C_{4}$ and $H_{2}\cong K_{1,q_{2}}$, where $4+q_{2}=q$.
  • (c)
    $H_{1}\cong H_{2}\cong C_{4}$ in the case of $q=8$.
By direct calculation, it is not hard to verify that $\lambda(G)<\lambda(T_{n,2,q})$. Next, we finish the proof in another way by using the double-eigenvector technique. We define $G^{\prime}$ to be the graph obtained from $G$ by deleting all edges of $H_{1}$ and $H_{2}$, and adding a copy $H\cong K_{1,q}$ into $S\backslash V(H_{1})$. To obtain a contradiction, it suffices to show $\lambda(G^{\prime})>\lambda(G)$. Recall that $x_{S}=\sum_{u\in S}x_{u}$ and $x_{T}=\sum_{u\in T}x_{u}$. Without loss of generality, we may assume that $x_{T}\geq x_{S}$. Let $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$ and $\mathbf{y}=\left(y_{1},y_{2},\ldots,y_{n}\right)^{\top}$ be the Perron vector of $G^{\prime}$. To show $\lambda^{\prime}>\lambda$, we shall prove
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0.$
Since $G$ is the graph with the maximum spectral radius, we have $\lambda\geq\lambda^{\prime}>\frac{n}{2}$. Observe that $H\cong K_{1,q}$ and $\sum_{v\in V(H)}d_{H}^{2}(v)=q^{2}+q$. For each $v\in S$, we have
$\lambda^{\prime}y_{v}=y_{T}+\sum_{u\in N_{H}(v)}x_{u}\geq y_{T}+d_{H}(v)\cdot\min_{u\in S}\{y_{u}\}.$
As $\lambda^{\prime}y_{u}\geq y_{T}$ for any $u\in S$, we get $\min_{u\in S}\{y_{u}\}\geq\frac{y_{T}}{\lambda^{\prime}}$. Using the double counting gives
$\sum_{\{u,v\}\in E(H)}\lambda^{\prime}(y_{u}+y_{v})$ $=$ $\sum_{v\in V(H)}d_{H}(v)\cdot\lambda^{\prime}y_{v}$
$\geq$ $\sum_{v\in V(H)}d_{H}(v)y_{T}+\sum_{v\in V(H)}d_{H}^{2}(v)\cdot\min_{u\in S}\{y_{u}\}$
$\geq$ $2qy_{T}+(q^{2}+q)\frac{y_{T}}{\lambda^{\prime}}.$
Note that $\lambda\geq\lambda^{\prime}$ and $V(H)\subseteq S\backslash V(H_{1})$ is an independent set in $G$. Then $N_{G}(v)=T$ and $\lambda x_{v}=x_{T}$ for each $v\in V(H)$. It follows that
$\sum_{\{u,v\}\in E(H)}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)=\sum_{\{u,v\}\in E(H)}\lambda^{\prime}\left(y_{u}+y_{v}\right)x_{T}\geq\left(2q+\frac{q^{2}+q}{\lambda}\right)y_{T}x_{T}.$ (6)
Since $H_{1}$ is a star or a $4$-cycle, we have $\sum_{v\in V\left(H_{1}\right)}d_{H_{1}}^{2}(v)\leq q_{1}^{2}+q_{1}$. Let $x_{v^{*}}=\max\{x_{u}:u\in S\}$. Note that $\lambda x_{v^{*}}=x_{T}+\sum_{u\in N_{H_{1}}(v^{*})}x_{u}\leq x_{T}+qx_{v^{*}}$. Then
$x_{v^{*}}\leq\frac{x_{T}}{\lambda-q}.$
Therefore, we have
$\sum_{\{u,v\}\in E(H_{1})}\lambda\left(x_{u}+x_{v}\right)$ $=$ $\sum_{v\in V(H_{1})}d_{H_{1}}(v)\cdot\lambda x_{v}\leq\sum_{v\in V(H_{1})}d_{H_{1}}(v)\bigl{(}x_{T}+d_{H_{1}}(v)x_{v^{*}}\bigr{)}$
$\leq$ $\sum_{v\in V(H_{1})}d_{H_{1}}(v)x_{T}+\sum_{v\in V(H_{1})}d_{H_{1}}^{2}(v)x_{v^{*}}$
$\leq$ $2q_{1}x_{T}+(q_{1}^{2}+q_{1})\frac{x_{T}}{\lambda-q}.$
Note that $V(H_{1})$ is an independent set of the new graph $G^{\prime}$. For every $v\in V(H_{1})$, we have $N_{G^{\prime}}(v)=T$ and $\lambda^{\prime}y_{v}=y_{T}$. Thus, we obtain
$\sum_{\{u,v\}\in E(H_{1})}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)=\sum_{\{u,v\}\in E(H_{1})}\lambda\left(x_{u}+x_{v}\right)\cdot y_{T}\leq\left(2q_{1}+\frac{q_{1}^{2}+q_{1}}{\lambda-q}\right)x_{T}y_{T}.$ (7)
Similarly, we can get $\max_{u\in T}x_{u}\leq\frac{x_{S}}{\lambda-q}$ and
$\sum_{\{u,v\}\in E(H_{2})}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)\leq\left(2q_{2}+\frac{q_{2}^{2}+q_{2}}{\lambda-q}\right)x_{S}y_{S}.$ (8)
Considering the above inequalities, we can obtain
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
$=$ $\sum_{\{u,v\}\in E(H)}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)-\sum_{\{u,v\}\in E\left(H_{1}\right)\cup E\left(H_{2}\right)}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)$
[…]
$=$ $\left(2q_{2}+\frac{q^{2}+q}{\lambda}-\frac{q_{1}^{2}+q_{1}}{\lambda-q}\right)x_{T}y_{T}-\left(2q_{2}+\frac{q_{2}^{2}+q_{2}}{\lambda-q}\right)x_{S}y_{S}.$
To show the right hand side is positive, we denote
$c_{1}:=2q_{2}+\frac{q^{2}+q}{\lambda}-\frac{q_{1}^{2}+q_{1}}{\lambda-q}$
and
$c_{2}:=2q_{2}+\frac{q_{2}^{2}+q_{2}}{\lambda-q}.$
By computation, we can check that $q_{1}q_{2}\geq q-1$ and
$c_{1}-c_{2}=\frac{2q_{1}q_{2}\lambda-q^{3}-q^{2}}{\lambda(\lambda-q)}>\frac{(q-1)n-q^{3}-q^{2}}{(n/2+q)n/2}>\frac{1}{10n},$
where we used the assumption ${n}/{2}<\lambda<{n}/{2}+q$ and $n\geq 300q^{2}$. Observe that $T$ is an independent set in $G^{\prime}$. Then for each $u\in T$, we have $\lambda^{\prime}y_{u}=y_{S}$. Summing over all $u\in T$ yields $|T|y_{S}=\lambda^{\prime}y_{T}=\lambda^{\prime}(y_{V}-y_{S})$. Thus we get
$y_{S}=\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}.$
Using Lemma 5.5 and $|S|=|T|=\frac{n}{2}$, it follows that
$y_{S}-y_{T}\leq\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}-\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|S|+\frac{2q}{\lambda^{\prime}-q}}<\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}.$
Combining with the assumption that $x_{T}\geq x_{S}$, we get
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
$\geq$ $c_{1}x_{T}y_{T}-c_{2}x_{S}y_{S}$
$\geq$ $\left(c_{2}+\frac{1}{10n}\right)x_{S}\left(y_{S}-\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}\right)-c_{2}x_{S}y_{S}$
$>$ $\frac{x_{S}}{10n}\cdot\frac{\lambda^{\prime}y_{V}}{n+q}-(2q-1)x_{S}\cdot\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V},$
where the last inequality holds by $y_{S}=\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}>\frac{\lambda^{\prime}y_{V}}{n+q}$ and $c_{2}<2(q-1)+\frac{q^{2}}{n/2-q}$. Hence, we can obtain $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0$ for $n\geq 300q^{2}$. This a contradiction with $\lambda^{\prime}>\lambda$. Thus, there exists exactly one $i\in\{1,2\}$ such that $H_{i}\neq\varnothing$ and then $H_{i}\cong K_{1,q}$. ∎
6 Proof of Theorem 2.5
Now, we are in a position to prove Theorem 2.5.
Proof of Theorem 2.5.
Assume that $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, where $s\in\mathbb{N}$. If $s=0$, then there is nothing to show. If $s=1$, then $G\neq T_{n,2}$. Theorem 1.3 implies $G$ has $\lfloor n/2\rfloor-1$ triangles. For the case $s=2$, it was proved in [27] that $G$ contains at least $n-3$ triangles. Now, we consider the case $s\geq 3$. We may assume that $t(G)\leq s{(n-1)}/{2}$. Otherwise, we are done.
Claim 1. We have $e(G)\geq\lfloor n^{2}/4\rfloor-3s$. There is a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6s$, $e(S,T)>\lfloor n^{2}/4\rfloor-9s$ and ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$.
Proof of Claim 1.
Since $\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>{(n-1)}/{2}$, by Lemma 3.2, we have
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-3s.$
If $G$ is $6s$-far from being bipartite, then using Theorem 3.3, we get
$t(G)\geq\frac{n}{6}\left(e(G)+6s-\frac{n^{2}}{4}\right)>s\frac{n}{2},$
which contradicts with the assumption. Thus, $G$ is not $6s$-far from being bipartite. Namely, there exists a vertex partition $V(G)=S\cup T$ such that
$e(S)+e(T)<6s.$
Consequently, we get $e(S,T)=e(G)-e(S)-e(T)>\lfloor{n^{2}}/{4}\rfloor-9s$. By the symmetry, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|<{n}/{2}-3\sqrt{s}$. Then $|T|=n-|S|>{n}/{2}+3\sqrt{s}$. It follows that $e(S,T)\leq|S||T|\leq\lfloor n^{2}/4\rfloor-9s$, which is a contradiction. Thus, we have ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$. ∎
Next, we are going to refine the above bipartition.
Claim 2. We have $e(S)+e(T)=s$.
Proof of Claim 2.
Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$. If $e(S)+e(T)\leq s-1$, then all triangles of $G$ can be covered by a vertex set of at most $s-1$ vertices, contradicting with $\tau_{3}(G)\geq s$. Now, suppose that $e(S)+e(T)\geq s+1$. Note that each edge of $G[S]$ or $G[T]$ can yield at least ${n}/{2}-3\sqrt{s}$ triangles in $K_{|S|,|T|}$. From the above discussion, we know that $e(S)+e(T)<6s$ and $|S||T|-e(S,T)\leq\lfloor n^{2}/4\rfloor-e(S,T)<9s$. Then $G[S,T]$ misses less than $9s$ edges between $S$ and $T$, and each missing edge destroys at most $6s$ triangles. So $G$ contains more than $(s+1)({n}/{2}-3\sqrt{s})-9s\cdot 6s>(s+1){n}/{2}-56.5s^{2}\geq s{n}/{2}$ triangles, where the last inequality holds for every $n\geq 113s^{2}$. This contradicts with the assumption $t(G)\leq s{(n-1)}/{2}$. Thus, we have $e(S)+e(T)=s$, as claimed. ∎
From the above claim, we obtain $e(S,T)=e(G)-s>\lfloor n^{2}/4\rfloor-4s$. Then
$\frac{n}{2}-2\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+2\sqrt{s}.$
Specifically, if $s=3$, then we have $\lceil n/2\rceil-3\leq|S|,|T|\leq\lfloor n/2\rfloor+3$. Since $e(S)+e(T)=s$ and there are less than $4s$ missing edges in $G[S,T]$, we know that $G$ contains at least $s(\lceil n/2\rceil-3)-4s^{2}\geq sn/2-5s^{2}$ triangles. If $s\geq 4$, then a similar argument shows that $G$ has at least $s({n}/{2}-2\sqrt{s})-4s^{2}\geq s{n}/{2}-5s^{2}$ triangles. ∎
7 Concluding remarks
Recall that $t(G)$ denotes the number of triangles in a graph $G$. A special case of an aforementioned result of Bollobás and Nikiforov [7] states that
$t(G)\geq\frac{n^{2}}{12}\left(\lambda-\frac{n}{2}\right).$
In fact, a careful examination indicates that the above inequality holds strictly. Motivated by this observation, one may ask the following problem.
Problem 7.1.
Finding the limit superior of $C$ such that
$t(G)\geq Cn^{2}\left(\lambda-\frac{n}{2}\right).$
Remark. By the clique density theorem of Reiher [49], we know that if $\varepsilon>0$ and $G$ is a graph on $n$ vertices with $e(G)\geq(\frac{1}{4}+\varepsilon)n^{2}$, then $G$ has at least $(\frac{1}{2}+o(1))\varepsilon n^{3}$ triangles, where $o(1)\to 0$ whenever $\varepsilon\to 0$. With aid of this powerful result, one can easily prove that if $G$ has at least $(\frac{1}{4}+\varepsilon)n^{2}$ edges, where we set $\varepsilon=\frac{1}{2}(\frac{\lambda}{n}-\frac{1}{2})$, then $G$ has at least $(\frac{1}{4}+o(1))n^{2}(\lambda-\frac{n}{2})$ triangles. Moreover, the graph $T_{n,2,1}$ shows that the coefficient $\frac{1}{4}$ is achievable. Very recently, Shengtong Zhang told us that taking $G=T_{n,3}$ yields $C\leq\frac{2}{9}$.
Another question as to how many triangles appear at least in any $n$-vertex graph with given spectral radius turns out to be interesting. For instance, for any graph $G$ with $\lambda(G)=\gamma n$, where $\gamma\in(\frac{1}{2},\frac{2}{3}]$, that is, $\lambda(T_{n,2})<\lambda(G)\leq\lambda(T_{n,3})$, there might exist a graph, say $L$, whose shape is approximately between $T_{n,2}$ and $T_{n,3}$, and $\lambda(L)=\gamma n$ such that $t(L)\leq t(G)$. Formally, we now define the minimizer $L$ as below. For $s\in\mathbb{N}$ and $\alpha>0$, we define $L_{n,s,\alpha}$ as a complete multipartite graph with $V(L_{n,s,\alpha})=V_{1}\cup\cdots V_{s}\cup V_{s+1}$, where $|V_{1}|=\cdots=|V_{s}|=\frac{n(1+\alpha)}{s+1}$ and $|V_{s+1}|=\frac{n(1-s\alpha)}{s+1}$. Next, we conjecture that $L_{n,s,\alpha}$ contains the minimum number of triangles among all graphs with given spectral radius.
Conjecture 7.2 (Spectral triangle density conjecture).
Let $\gamma\in(\frac{1}{2},1)$ and $G$ be an $n$-vertex graph with $\lambda(G)\geq\gamma n$. Then $t(G)\geq t(L_{n,s,\alpha})$, where $s\geq 2$ is an integer with $\gamma\in(\frac{s-1}{s},\frac{s}{s+1}]$ and $\alpha>0$ is a maximal number such that $\lambda(L_{n,s,\alpha})\geq\gamma n$.
Nosal [47] proved that every graph $G$ with $m$ edges and $\lambda(G)>\sqrt{m}$ contains a triangle; see, e.g., [38, 44]. Recently, Ning and Zhai [45] showed that if $\lambda(G)\geq\sqrt{m}$, then $t(G)\geq\lfloor\frac{\sqrt{m}-1}{2}\rfloor$, unless $G$ is a complete bipartite graph. We propose the following conjecture.
Conjecture 7.3.
If $G$ is a graph with $m$ edges and
$\lambda(G)\geq\frac{1+\sqrt{4m-3}}{2},$
then $G$ has at least $\frac{m-1}{2}$ triangles, with equality if and only if $G=K_{2}\vee\frac{m-1}{2}K_{1}$.
Remark. Roughly speaking, Conjecture 7.3 reveals that in a graph $G$ with $\lambda(G)\approx\sqrt{m}+\frac{1}{2}$, the number of triangles jumps from $\frac{1}{2}\sqrt{m}$ to $\frac{1}{2}m$. In addition, we can see that if $\lambda(G)\geq(1+\varepsilon)\sqrt{m}$, then $t(G)>\frac{2\varepsilon}{3}m^{3/2}$. In fact, it is not difficult to obtain a weaker bound than Conjecture 7.3. Indeed, it follows from Lemma 3.2 that $t(G)>\frac{1}{3}\lambda(\lambda^{2}-m)\geq\frac{m-1}{3}$. Thus, the difficulty of Conjecture 7.3 lies in deriving the tight bound $t(G)\geq\frac{m-1}{2}$.
Recall in Theorem 1.2 that for every $q<n/2$, the graph $G$ with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles. While, in our Theorem 2.2, we proved the spectral version for $n\geq 300q^{2}$. Under our framework, it seems possible to improve this bound slightly. For the sake of clarity, we do not make any serious attempt to optimize the absolute constant in our statements and proofs. It is ambitious to extend our result to the following wide range.
Problem 7.4.
Whether Theorems 2.2 and 2.3 hold for every $q<n/2$? If not, what is the limit superior of $\delta$ such that the spectral results hold for $q<\delta n$?
In 2017, Pikhurko and Yilma [48, Theorem 1.4] proved that for every color-critical graph $F$ with $\chi(F)=r+1$, there exist $\delta=\delta(F)>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and $q<\delta n$, if $G$ is an $n$-vertex graph with $e(T_{n,r})+q$ edges and it has the minimum number of copies of $F$, then $G$ is obtained from $T_{n,r}$ by adding $q$ new edges. Theorem 1.2 gives $\delta(K_{3})=1/2$. More generally, a result of Lovász and Simonovits [34, 35] implies that $\delta(K_{r+1})=1/r$ for every $r\geq 2$. Moreover, Pikhurko and Yilma [48] also determined the limit superior of $\delta$ for a number of color-critical graphs, such as, odd cycles $\delta(C_{2k+1})=1/2$, and cliques with one edge removed $\delta(K_{r+2}-e)=(r-1)/r^{2}$. In a forthcoming paper, we will use a variety of different methods to establish a spectral counting result for any color-critical graph. As the techniques are quite complicated, we intend to present it separately $\heartsuit$.
Declaration of competing interest
The authors declare that they have no conflicts of interest to this work.
Data availability
No data was used for the research described in the article.
Acknowledgements
The authors are thankful to Xiaocong He and Loujun Yu for carefully reading an early manuscript of this paper. Yongtao Li would like to thank Dheer Noal Desai for the inspiring discussions on spectral graph problems over the years. Li was supported by the Postdoctoral Fellowship Program of CPSF (No. GZC20233196), Feng was supported by the NSFC grant (No. 12271527), and Peng was supported by the NSFC grant (No. 11931002).

### v2

@@
-Yongtao Li, Lihua Feng, Yuejian Peng
+A spectral Lovász–Simonovits theorem
 Abstract
-A fundamental result in extremal graph theory attributes to Mantel’s theorem, which states that every graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle.
-About half of a century ago, Lovász and Simonovits (1975) provided a supersaturation phenomenon, which asserts that for $q<n/2$, every graph with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles.
-This result solved a conjecture proposed by Erdős in 1962.
-In this paper, we establish a spectral version of the result of Lovász and Simonovits.
-Let $Y_{n,2,q}$ be the graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding a matching with $q$ edges into the vertex part of size $\lceil n/2\rceil$.
-Using the supersaturation-stability method and the classical spectral techniques, we firstly prove that for $n\geq 300q^{2}$, each graph $G$ on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$ contains at least $q\lfloor n/2\rfloor$ triangles.
-Moreover, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star with $q$ edges into the vertex part of size $\lceil n/2\rceil$.
-Secondly, we show further that $T_{n,2,q}$ is the unique spectral extremal graph that contains at most $q\lfloor n/2\rfloor$ triangles and attains the maximum of the spectral radius.
-This result answers a spectral triangle counting problem due to Ning and Zhai (2023).
-Thirdly, we present an asymptotically spectral stability result under a specific constraint on the triangle covering number.
-The third result could be regarded as a spectral extension of a recent result proved by Balogh and Clemen (2023), and independently by Liu and Mubayi (2022).
-Key words: Extremal graph theory; counting triangles; adjacency matrix.
+A fundamental result in extremal graph theory is attributed to Mantel’s theorem, which states that every graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges must contain a triangle.
+Lovász and Simonovits (1975) provided a supersaturation phenomenon by showing that for any $q<n/2$, every graph with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles.
+This result resolved a conjecture proposed by Erdős in 1962.
+In this paper, we establish a spectral counterpart of the result of Lovász and Simonovits.
+Let $Y_{n,2,q}$ be the graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding a matching with $q$ edges into the partite set of size $\lceil n/2\rceil$.
+Using the supersaturation-stability method and the spectral techniques, we firstly prove that for $q\leq\frac{1}{11}\sqrt{n}$, every graph $G$ on $n$ vertices with spectral radius $\lambda(G)\geq\lambda(Y_{n,2,q})$ contains at least $q\lfloor n/2\rfloor$ triangles.
+We also show that the bound $q=O(\sqrt{n})$ is tight up to a constant factor, yielding a phenomenon different from that in edge supersaturation.
+Our result answers a spectral triangle counting problem proposed by Ning and Zhai (2023).
+Secondly, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star with $q$ edges into the partite set of size $\lceil n/2\rceil$.
+We show further that $T_{n,2,q}$ is the unique extremal graph that contains at most $q\lfloor n/2\rfloor$ triangles and attains the maximum spectral radius.
+Thirdly, we present an asymptotic spectral stability result under a specific constraint on the triangle covering number.
+This result could be viewed as a spectral extension of a recent result proved by Balogh and Clemen (2023), and independently by Liu and Mubayi (2022).
+Key words: Spectral graph theory; counting triangles; adjacency matrix.
 2010 Mathematics Subject Classification. 05C35; 05C50.
@@
 This extremal number is denoted by $\mathrm{ex}(n,F)$.
-It is a longstanding open problem in extremal combinatorics to develop some understanding on the function $\mathrm{ex}(n,F)$ for a general graph $F$.
+It is a longstanding open problem in extremal combinatorics to develop some understanding of the function $\mathrm{ex}(n,F)$ for a general graph $F$.
 A well-known result of Mantel (see [4, 36]) asserts that an $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle.
@@
 Shortly afterwards, there have been various generalizations of Mantel’s theorem; see, e.g., [5, 6].
-1.1 Extremal graph problems on triangles
-About seventy years ago, Rademacher (unpublished, see Erdős [15, 18]) showed an interesting extension on Mantel’s theorem.
-Namely, there are not just one but at least $\lfloor{n}/{2}\rfloor$ triangles in any $n$-vertex graph with more than $\lfloor{n^{2}}/{4}\rfloor$ edges.
-Theorem 1.1 (Erdős–Rademacher, 1955).
-If $G$ is a graph on $n$ vertices with
-$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+1,$
-then $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G$ is the graph obtained from $T_{n,2}$ by adding one edge into the vertex part of size $\lceil{n}/{2}\rceil$.
-We remark that Erdős [18] proved a variant for a graph with less than $\lfloor n^{2}/4\rfloor$ edges.
-The result of Erdős states that if $\ell\geq 0$ and $G$ is an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor-\ell$, and $G$ has a triangle, then it has at least $\lfloor{n}/{2}\rfloor-\ell-1$ triangles.
-As a special case, we can see that if $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$, then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G=T_{n,2}$.
-The problem of determining the minimum number of triangles is usually referred to as the Erdős–Rademacher problem, which is regarded as the starting point of the supersaturation problem in extremal graph theory.
-After proving Theorem 1.1, Erdős [16, 17] showed that there exists a small constant $c>0$ such that if $1\leq q<cn$ and $n$ is sufficiently large, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles.
-This result was further generalized by Lovász and Simonovits [34, 35] as follows.
+1.1 Extremal graph results on triangles
+About eighty years ago, Rademacher (unpublished, see Erdős [15, 18]) provided an interesting extension on Mantel’s theorem by showing that there are not just one but at least $\lfloor{n}/{2}\rfloor$ triangles in any $n$-vertex graph with more than $\lfloor{n^{2}}/{4}\rfloor$ edges.
+Theorem 1.1 (Erdős–Rademacher, 1941).
+If $G$ is a graph on $n$ vertices with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor+1$, then $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G$ is the graph obtained from $T_{n,2}$ by adding one edge into the partite set of size $\lceil{n}/{2}\rceil$.
+The problem of counting triangles is referred to as the Erdős–Rademacher problem, which is regarded as a starting point of the supersaturation in graph theory.
+Erdős [18] proved a variant for a graph with less than $\lfloor n^{2}/4\rfloor$ edges.
+The Erdős result states that if $\ell\geq 0$ and $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$-$\ell$, and $G$ has a triangle, then it has at least $\lfloor{n}/{2}\rfloor-\ell-1$ triangles.
+Setting $\ell=0$, we see that if $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$, then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G=T_{n,2}$.
+Moreover, Erdős [16, 17] showed that there exists $c>0$ such that if $1\leq q<cn$ and $n$ is sufficiently large, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles.
+This was further generalized by Lovász and Simonovits [37, 38].
 Theorem 1.2 (Lovász–Simonovits, 1975).
-Let $1\leq q<{n}/{2}$ be an integer and $G$ be an $n$-vertex graph with
-$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+q.$
-Then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles.
+Let $1\leq q<{n}/{2}$ be an integer and $G$ be an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor+q$. Then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles.
 A quick proof of Theorem 1.2 in the range of $q\leq{n}/{4}$ can be found in [4, page 302].
-Incidentally, we mention that the condition $q<{n}/{2}$ is necessary.
-For even $n$ and $q={n}/{2}$, one can add $q+1$ extra edges to the larger vertex part of $K_{\frac{n}{2}+1,\frac{n}{2}-1}$ such that there is no triangle among these $q+1$ edges.
-Hence, we get a graph with ${n^{2}}/{4}+q$ edges and $(q+1)({n}/{2}-1)={n^{2}}/{4}-1$ triangles, which is less than $q\lfloor{n}/{2}\rfloor$ triangles.
+Incidentally, we mention that the condition q<n/2q<{n}/{2} is necessary.
+For even $n$ and q=n/2q={n}/{2}, one can add $q+1$ extra edges to the larger vertex part of $K_{\frac{n}{2}+1,\frac{n}{2}-1}$ such that there is no triangle among these $q+1$ edges.
+Hence, we get a graph with ${n^{2}}/{4}+q$ edges and $(q+1)({n}/{2}-1)={n^{2}}/{4}-1$ triangles, which are less than $q\lfloor{n}/{2}\rfloor$ triangles.
 The stabilities of Theorems 1.1 and 1.2 were recently studied by Xiao and Katona [59], Liu and Mubayi [34] and Balogh and Clemen [3].
 A graph is called color-critical if it has an edge whose deletion reduces its chromatic number.
-For example, the complete graph, odd cycle, book and even wheel are color-critical.
+For example, the complete graphs and odd cycles are color-critical.
 Apart from the triangles, the supersaturation problem for color-critical graphs was also investigated by Mubayi [40], Pikhurko and Yilma [51], Ma and Yuan [39].
@@
 1.2 Eigenvalues and counting triangles
-Spectral graph theory aims to apply the eigenvalues of associated matrices to investigate the structural information of a graph.
-Let $G$ be a simple graph on $[n]$.
-The adjacency matrix is defined as $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $i$ and $j$ are adjacent, and $a_{i,j}=0$ otherwise.
-Let $\lambda(G)$ be the spectral radius of $G$, which is defined as the maximum of modulus of eigenvalues of $A(G)$.
-By the Perron–Frobenius theorem, we know that $\lambda(G)$ is the largest eigenvalue of $A(G)$ and there exists a nonnegative eigenvector corresponding to $\lambda(G)$.
-As mentioned before, the Turán type problem studies the condition on the number of edges of a graph to force certain subgraphs, e.g., the classical Turán theorem states that $e(G)>(1-\nicefrac{{1}}{{r}})n^{2}/2$ forces a clique $K_{r+1}$.
-Correspondingly, one could wish to investigate the condition on spectral radius of the associated adjacency matrix to guarantee certain subgraphs.
-The interplay between these two areas above is called the spectral Turán type problem.
-Finding general conditions on the spectral radius which ensure that $G$ has specific substructures or properties is one of the most popular topic in Spectral Extremal Graph Theory.
-One of the well-known results on this topic was obtained in 1986 by Wilf [54] who showed that every $n$-vertex graph $G$ with $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$ has a clique $K_{r+1}$.
-Another significant result due to Nikiforov [38] states that an $m$-edge graph $G$ with $\lambda^{2}(G)>{(1-\nicefrac{{1}}{{r}})2m}$ contains a clique $K_{r+1}$.
-The above spectral results generalize the Turán theorem.
-As illustrated by Nikiforov, his result can imply Wilf’s result by assuming Turán’s theorem.
-Actually, this implication can also be seen by the fact $\lambda(G)\geq 2m/n$.
-Indeed, we know that if $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$, then $\lambda^{2}(G)>(1-\nicefrac{{1}}{{r}})n\cdot(2m/n)=(1-\nicefrac{{1}}{{r}})2m$, as expected.
-In addition, the spectral radius of a graph is always at least the minimum degree, if $G$ is a $K_{r+1}$-free graph on $n$ vertices, then Wilf’s result implies $\delta(G)\leq(1-\nicefrac{{1}}{{r}})n$ and so $\delta(G)\leq n-\lceil n/r\rceil=\delta(T_{n,r})$.
-Taking a vertex $u\in V(G)$ with degree $d(u)=\delta(G)$, it follows by induction that $e(G)=e(G\setminus\{u\})+\delta(G)\leq e(T_{n-1,r})+\delta(T_{n,r})=e(T_{n,r})$.
-We conclude that Wilf’s result can imply the “strong” Turán theorem as well.
-The spectral Turán problems have been receiving considerable attention in the last two decades and it is an important topic in modern graph theory; see, e.g., [38, 7, 23] for cliques, [20] for matchings, [12] for even cycles, [21, 62, 22] for cycles of consecutive lengths, [39] for color-critical graphs, [1, 42] for complete bipartite graphs, [51, 28] for planar and outerplanar graphs, [52, 59, 8, 61] for graphs without certain minors, [40] for a spectral Erdős–Stone–Bollobás theorem, [41] for the spectral stability theorem, [11] for a spectral Erdős–Sós theorem, [57] for a spectral Erdős–Pósa theorem, [25] for a spectral Erdős–Rademacher theorem, [26] for a spectral Erdős–Faudree–Rousseau theorem, [9, 58] for friendship graphs, [24, 27] for bowties, [60, 43] for book graphs, [10] for odd wheels and [53, 19] for a characterization of spectral extremal problems for a class of graphs.
-Although there has been a wealth of research on the spectral extremal problems in recent years, there are very few spectral results on the problem of counting substructures.
-In particular, Bollobás and Nikiforov [7] proved that every $n$-vertex graph has at least $(\frac{\lambda}{n}-1+\frac{1}{r})\frac{r(r-1)}{r+1}(\frac{n}{r})^{r+1}$ copies of $K_{r+1}$.
-This result provides a spectral supersaturation on cliques.
-Moreover, the spectral supersaturation for a general graph can be found in [25].
-Let $F$ be a graph with chromatic number $\chi(F)=r+1$.
-For any $\varepsilon>0$, there exist $\delta>0$ and $n_{0}\in\mathbb{N}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with $\lambda(G)\geq\left(1-\frac{1}{r}+\varepsilon\right)n$, then $G$ contains at least $\delta n^{f}$ copies of $F$, where $f$ is the number of vertices of $F$.
-Recall in Theorem 1.1 that $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles whenever $e(G)>e(T_{n,2})$.
-Correspondingly, it is natural to consider the spectral version: if $G$ is a graph with $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles.
+Spectral graph theory aims to apply algebraic properties, such as the eigenvalues of associated matrices, to investigate the structural properties of a graph.
+In recent years, significant breakthroughs have been made in understanding graph structure through matrix eigenvalues; see, e.g., [21, 22].
+Let $G$ be a graph on $n$ vertices, denoted by $V(G)=\{v_{1},\ldots,v_{n}\}$.
+The adjacency matrix of $G$ is denoted by $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $v_{i}$ and $v_{j}$ are adjacent, and $a_{i,j}=0$ otherwise.
+The spectral radius $\lambda(G)$ is defined as the maximum modulus of the eigenvalues of $A(G)$.
+By the Perron–Frobenius theorem, $\lambda(G)$ is the largest eigenvalue of $A(G)$ and there exists a non-negative eigenvector corresponding to it.
+The Rayleigh quotient gives $\lambda(G)=\max_{\bm{x}\in\mathbb{R}^{n}}\frac{\bm{x}^{\mathrm{T}}A(G)\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$, which implies $\lambda(G)\geq\frac{2m}{n}\geq\delta(G)$, where each equality holds if and only if $G$ is a regular graph.
+The classical Turán theorem states that $e(G)>(1-\nicefrac{{1}}{{r}})n^{2}/2$ forces a clique $K_{r+1}$.
+Similarly, one may wish to investigate the condition on spectral radius $\lambda(G)$ to guarantee certain substructures.
+Such problems fall under the umbrella of spectral extremal graph theory, which is one of the most popular topic in spectral graph theory.
+A well-known result on this topic was obtained in 1986 by Wilf [57], who showed that every $n$-vertex graph $G$ with $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$ has a clique $K_{r+1}$.
+Another significant result due to Nikiforov [41] states that every $m$-edge graph $G$ with $\lambda^{2}(G)>{(1-\nicefrac{{1}}{{r}})2m}$ contains a clique $K_{r+1}$.
+The above spectral results generalize the aforementioned Turán theorem.
+Nikiforov’s result implies Wilf’s result by invoking Turán’s bound or Rayleigh’s bound $\lambda(G)\geq 2m/n$.
+Indeed, if $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$, then $\lambda^{2}(G)>(1-\nicefrac{{1}}{{r}})n\cdot(2m/n)=(1-\nicefrac{{1}}{{r}})2m$, as desired.
+In addition, Wilf’s bound implies the “strong” Turán theorem.
+To see this, if $G$ is a $K_{r+1}$-free graph, then $\delta(G)\leq\lambda(G)\leq(1-\nicefrac{{1}}{{r}})n$ by assuming Wilf’s bound.
+Hence $\delta(G)\leq n-\lceil n/r\rceil=\delta(T_{n,r})$.
+Taking a vertex $u\in V(G)$ with degree $d(u)=\delta(G)$.
+By induction, we get $e(G\setminus\{u\})\leq e(T_{n-1,r})$, which yields $e(G)=e(G\setminus\{u\})+\delta(G)\leq e(T_{n-1,r})+\delta(T_{n,r})=e(T_{n,r})$, as expected.
+The spectral Turán problems have been receiving considerable attention in the last two decades and it is an important topic in modern graph theory; see, e.g., [41, 7, 25] for cliques, [20] for matchings, [12] for even cycles, [23, 65, 24] for cycles of consecutive lengths, [42] for color-critical graphs, [1, 45] for complete bipartite graphs, [54, 30] for planar and outerplanar graphs, [55, 62, 8, 66] for graphs without certain minors, [43] for a spectral Erdős–Stone–Bollobás theorem, [44] for the spectral stability theorem, [11] for a spectral Erdős–Sós theorem, [60] for a spectral Erdős–Pósa theorem, [27] for a spectral Erdős–Rademacher theorem, [28] for a spectral Erdős–Faudree–Rousseau theorem, [9, 61, 28] for friendship graphs, [26, 29] for bowties, [63, 46] for book graphs, [10] for odd wheels and [56, 19] for a characterization of spectral-consistent graphs.
+Although there has been a wealth of research on the spectral extremal problems, there are very few results on the spectral problem of counting substructures.
+In 2007, Bollobás and Nikiforov [7] proved that every $n$-vertex graph $G$ has at least $\big(\frac{\lambda(G)}{n}-1+\frac{1}{r}\big)\frac{r(r-1)}{r+1}(\frac{n}{r})^{r+1}$ copies of $K_{r+1}$.
+This result provides a spectral supersaturation for cliques.
+The spectral supersaturation for general graphs (see [27, Sec.
+5]) states that for any $\varepsilon>0$ and any graph $F$ with $\chi(F)=r+1$, there exist $\delta>0$ and $n_{0}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with $\lambda(G)\geq\left(1-\frac{1}{r}+\varepsilon\right)n$, then $G$ contains at least $\delta n^{|F|}$ copies of $F$.
+Recall in Theorem 1.1 that $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles when $e(G)>e(T_{n,2})$.
+Correspondingly, it is natural to consider the spectral version: if $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles.
 However, this intuition is not true.
-Let $K_{a,b}^{+}$ be the graph obtained from the complete bipartite graph $K_{a,b}$ by adding an edge to the vertex set of size $a$.
-For even $n$, we take $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$.
-It is easy to verify that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles.
-Fortunately, it was proved by Ning and Zhai [45] that $\lfloor\frac{n}{2}\rfloor-1$ triangles can be guaranteed, which establishes a spectral version of a result of Erdős [18]; see the remark after Theorem 1.1.
+Let $K_{a,b}^{+}$ be the graph obtained from $K_{a,b}$ by adding an edge to the vertex set of size $a$.
+For even $n$, taking $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$, we compute that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles.
+Ning and Zhai [48] proved that $\lfloor{n}/{2}\rfloor-1$ triangles can be guaranteed.
 Theorem 1.3 (Ning–Zhai, 2023).
@@
 then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G$ is the bipartite Turán graph $T_{n,2}$.
-Furthermore, Ning and Zhai [45] proposed a general problem on finding spectral counting results for color-critical graphs, including, e.g., the spectral version of the Lovász–Simonovits result in Theorem 1.2.
-Incidentally, this problem was also selected as one of the “Unsolved Problems” in spectral graph theory [31, Section 9].
-To the best of our knowledge, there are few spectral results on counting substructures in a graph with given order, including only triangles [45], cliques [7], triangular edges [26] and bowties [27].
+Ning and Zhai [48] proposed a problem on finding spectral counting results for color-critical graphs, including, e.g., the spectral version of Theorem 1.2.
+Interestingly, this problem was also selected as one of the “Unsolved Problems” in spectral graph theory [33, Sec.
+9].
+There are few spectral results on counting substructures in a graph with given order, including only triangles [48], cliques [7], triangular edges [28] and bowties [29].
 The main contribution of this paper is to establish a spectral version of Theorem 1.2.
-To some extent, our result (in next section) could be viewed as a partial resolution of the Ning–Zhai problem.
+To some extent, our result could be viewed as a partial resolution of the Ning–Zhai problem.
 2 Main results
 2.1 A spectral Lovász–Simonovits theorem
-Before showing a spectral version of Theorem 1.2, the first obstacle one may ask is to establish a spectral condition corresponding to the conventional size-condition.
-How to determine an increment function $\delta(q)>0$ such that the condition $e(G)\geq\lfloor n^{2}/4\rfloor+q$ can be replaced with the spectral condition $\lambda(G)\geq\lambda(T_{n,2})+\delta(q)$.
-For instance, is it sufficient to take $\delta(q)=2q/n$?
-As seen from the phenomenon, where every additional edge guarantees $\lfloor n/2\rfloor$ additional triangles, it is unlikely that such an increment on the spectral radius can create $\lfloor n/2\rfloor$ new triangles.
+Before trying to obtain the spectral counterpart of Theorem 1.2, the first obstacle we may encounter is how to formulate a spectral hypothesis that corresponds to the conventional assumption on the size of a graph.
+More precisely, we would like to determine an increment function $\delta(q)>0$ such that the condition $e(G)\geq\lfloor n^{2}/4\rfloor+q$ can be replaced with a spectral condition of the form $\lambda(G)\geq\lambda(T_{n,2})+\delta(q)$.
+For instance, is it sufficient to consider $\delta(q)=2q/n$?
+Since every additional edge guarantees $\lfloor n/2\rfloor$ additional triangles, it is unlikely that such spectral increment can create $\lfloor n/2\rfloor$ new triangles.
 From this point of view, it seems difficult to measure the increment on the spectral radius of a graph.
-To get around this difficulty, Li, Lu and Peng [25] presented a new way to measure the spectral increment.
-To start with, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the larger vertex part; see Figure 1.
-Under this notation, they define the spectral increment as $\delta(q)=\lambda(T_{n,2,q})-\lambda(T_{n,2})$.
-In particular, they [25, Problem 5.3] proposed the following problem.
+To get around the difficulty, Li, Lu and Peng [27] presented a new way to measure the increment on the spectral radius.
+Let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the larger vertex part; see Fig.
+1.
+Under this notation, they define the spectral increment as $\delta(q)=\lambda(T_{n,2,q})-\lambda(T_{n,2})$, and proposed the following problem.
 Problem 2.1 (Li–Lu–Peng, 2024).
@@
 Then $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G=T_{n,2,q}$.
-Problem 2.1 could be regarded as a spectral version of the Lovász–Simonovits theorem.
-Especially, Li, Lu and Peng [25] confirmed Problem 2.1 in the fundamental case of $q=1$.
-In the present paper, our goal is to solve this problem for a general $q\in\mathbb{N}$.
+Problem 2.1 could be regarded as a spectral version of the Lovász–Simonovits theorem. Especially, Li, Lu and Peng [27] confirmed this problem in the initial case of $q=1$.
+In the present paper, our goal is to confirm Problem 2.1 for any given $q\in\mathbb{N}$.
 Let $t(G)$ be the number of triangles in $G$.
-Moreover, we will show further that $t(G)\geq q\lfloor{n}/{2}\rfloor$ in Problem 2.1 can be guaranteed by a slightly weaker condition.
-Let $Y_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the vertex part of larger size; see Figure 1.
-Clearly, we can see that $e(Y_{n,2,q})=e(T_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ and $t(Y_{n,2,q})=t(T_{n,2,q})=q\lfloor n/2\rfloor$.
-Upon direct computation, it is not difficult to verify that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
+We show that $t(G)\geq q\lfloor{n}/{2}\rfloor$ in Problem 2.1 can be guaranteed by a slightly weaker condition.
+Let $Y_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the partite set with larger size; see Fig.
+1.
+It is not difficult to verify that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$ for every $q\geq 2$; see Proposition 3.2.
 Figure 1: The graphs $Y_{n,2,q}$ and $T_{n,2,q}$.
-In what follows, we are going to present the spectral version of the Lovász–Simonovits theorem. First of all, we give a result stronger than Problem 2.1.
+In what follows, we give a stronger result than Problem 2.1.
 Theorem 2.2.
-If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with
+If $1\leq q\leq\frac{1}{11}\sqrt{n}$ and $G$ is an $n$-vertex graph with
 $\lambda(G)\geq\lambda(Y_{n,2,q}),$
@@
 Remark.
-The spectral condition in Theorem 2.2 is easier to fulfill than the size condition in Theorem 1.2.
-Indeed, invoking the fact $\lambda(G)\geq 2e(G)/n$, we can verify that if $e(G)>e(Y_{n,2,q})$, then $\lambda(G)>\lambda(Y_{n,2,q})$.
-Otherwise, assuming that $\lambda(G)\leq\lambda(Y_{n,2,q})$, it yields that $e(G)\leq\lfloor\frac{n}{2}\lambda(G)\rfloor\leq\lfloor\frac{n}{2}\lambda(Y_{n,2,q})\rfloor=e(Y_{n,2,q})$, where the last equality holds by $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{8q}{n^{2}}$.
-In conclusion, the spectral condition is weaker than the size condition.
-Additionally, there are a vast number of graphs $G$ with $\lambda(G)>\lambda(Y_{n,2,q})$ while $e(G)<e(Y_{n,2,q})$.
-Note that $\lambda(T_{n,2,q})>\lambda(Y_{n,2,q})$.
+The above bound $q\leq\frac{1}{11}\sqrt{n}$ is tight up to a constant factor; see Example 3.3.
+This reveals a different phenomenon from that of the traditional edge supersaturation.
+Moreover, the spectral condition in Theorem 2.2 is weaker than the size condition in Theorem 1.2, since $e(G)>e(Y_{n,2,q})$ implies $\lambda(G)>\lambda(Y_{n,2,q})$.
+Indeed, suppose that $\lambda(G)\leq\lambda(Y_{n,2,q})$.
+Then, invoking the fact that $\lambda(G)\geq\frac{2e(G)}{n}$, we obtain $e(G)\leq\lfloor\frac{n}{2}\lambda(G)\rfloor\leq\lfloor\frac{n}{2}\lambda(Y_{n,2,q})\rfloor=e(Y_{n,2,q})$, where the last equality follows from $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ (see Lemma 3.1).
+In conclusion, the spectral condition is indeed weaker than the size condition.
+In addition, there are many graphs $G$ satisfying $\lambda(G)>\lambda(Y_{n,2,q})$ while $e(G)<e(Y_{n,2,q})$.
+From this perspective, the spectral condition of Theorem 2.2 has wider applicability.
+The Bollobás–Nikiforov bound [7] states that $t(G)\geq\frac{n^{2}}{12}\left(\lambda(G)-\frac{n}{2}\right)$.
+This provides a lower bound for $t(G)$ under the assumption of Theorem 2.2.
+For even $n$, we have $\lambda(G)\geq\lambda(Y_{n,2,q})>\frac{n}{2}+\frac{2q}{n}$, and then $t(G)>\frac{q}{6}n$; For odd $n$, we obtain $\lambda(G)>\frac{n}{2}-\frac{1}{2n}+\frac{2q}{n}$, which implies $t(G)>\frac{4q-1}{24}n$.
+Since $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ for all $n\geq 4q$ (see Lemma 3.1), the lower bound on $t(G)$ provided by Bollobás–Nikiforov’s bound does not exceed $\frac{q}{6}n+\frac{q}{3}$.
+In contrast, Theorem 2.2 offers a stronger bound $t(G)\geq q\lfloor\frac{n}{2}\rfloor$.
+Recall that $\lambda(T_{n,2,q})>\lambda(Y_{n,2,q})$.
 Applying Theorem 2.2, we know that the desired bound in Problem 2.1 holds immediately.
-By using the double-eigenvector technique, we shall confirm Problem 2.1 by proving that $T_{n,2,q}$ is the unique graph that attains the maximum spectral radius among all $n$-vertex graphs with at most $q\lfloor{n}/{2}\rfloor$ triangles.
+Using the double-eigenvector technique, we shall confirm Problem 2.1 in the range $q\leq\frac{1}{11}\sqrt{n}$, showing that $T_{n,2,q}$ is the unique graph that attains the maximum spectral radius among all $n$-vertex graphs with at most $q\lfloor{n}/{2}\rfloor$ triangles.
 Theorem 2.3.
-If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with
+If $1\leq q\leq\frac{1}{11}\sqrt{n}$ and $G$ is an $n$-vertex graph with
 $\lambda(G)\geq\lambda(T_{n,2,q}),$
 then $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G=T_{n,2,q}$.
-When we embed a graph $H$ with $q$ edges into a vertex part of the bipartite Turán graph $T_{n,2}$, we can compute that the spectral radius of the resulting graph will be decreasing in the order of star, complete graph, complete bipartite graph, cycle, path and matching.
-In addition, we can verify that if we embed two matchings of size $q_{1}$ and $q_{2}$ with $q_{1}+q_{2}=q$ into two vertex parts, respectively, then the resulting graph has larger spectral radius than that of $Y_{n,2,q}$.
-On the contrary, if we embed two stars of sizes $q_{1}$ and $q_{2}$ into two parts, respectively, then the resulting graph has spectral radius smaller than that of $T_{n,2,q}$.
-2.2 A spectral extension on a stability result
-In 2021, Xiao and Katona [56] initiated the study on stability of Theorem 1.1.
-The triangle covering number $\tau_{3}(G)$ is defined as the minimum number of vertices in a set that contains at least one vertex of each triangle in $G$.
+Remark.
+When embedding a graph $H$ with $q$ edges into one partite set of $T_{n,2}$, we can verify that the spectral radius of the resulting graph decreases according to the following order of $H$: star, complete graph, complete bipartite graph (that is not a star), cycle, path, and matching.
+Moreover, it can be verified that if we embed two matchings of sizes $q_{1}$ and $q_{2}$ (with $q_{1}+q_{2}=q$) into two partite sets, respectively, then the resulting graph has larger spectral radius than $Y_{n,2,q}$.
+In contrast, if we embed two stars of sizes $q_{1}$ and $q_{2}$ into two partite sets, respectively, then the resulting graph has smaller spectral radius than $T_{n,2,q}$.
+2.2 A spectral extension of a stability result
+The triangle covering number $\tau_{3}(G)$ is the minimum number of vertices in a set that contains at least one vertex of each triangle in $G$.
 In particular, $\tau_{3}(G)=1$ if and only if all triangles of $G$ share a common vertex.
-It was shown in [56] that if $G$ has more than $\lfloor n^{2}/4\rfloor$ edges and $\tau_{3}(G)\geq 2$, then $G$ has at least $n-2$ triangles.
-Furthermore, they proposed a problem to study the stability of Theorem 1.2.
+Xiao and Katona [59] gave a stability of Theorem 1.1 by showing that if $e(G)>\lfloor n^{2}/4\rfloor$ and $\tau_{3}(G)\geq 2$, then $G$ has at least n2n-2 triangles.
+Moreover, they proposed a problem to study the stability of Theorem 1.2.
 Recently, this problem was solved by Liu and Mubayi [34], as well as Balogh and Clemen [3] independently.
@@
 Firstly, let $G_{1}$ be the graph of order $n$, where $V(G_{1})=A\cup B$ with $|A|=\lceil{n}/{2}\rceil+a$ and $|B|=\lfloor{n}/{2}\rfloor-a$, where $a\in\mathbb{N}$ is chosen later.
-Pick $2(s-1)$ vertices $x_{1},y_{1},\ldots,x_{s-1},y_{s-1}$ in $A$ and two vertices $u_{1},u_{2}$ in $B$.
-Adding $s$ edges $\{u_{1},u_{2}\}$ and $\{x_{i},y_{i}\}$ into $K_{|A|,|B|}$, and delete $\alpha$ edges $\{u_{1},x_{1}\},\ldots,\{u_{1},x_{\alpha}\}$, where $\alpha:=s-t-a^{2}-\mathbf{1}_{\{2\nmid n\}}a$.
+Pick 2(s1)2(s-1) vertices $x_{1},y_{1},\ldots,x_{s-1},y_{s-1}$ in $A$ and two vertices $u_{1},u_{2}$ in $B$.
+Adding $s$ edges $\{u_{1},u_{2}\}$ and $\{x_{i},y_{i}\}$, for $i\in\{1,\ldots,s-1\}$ into $K_{|A|,|B|}$, and delete $\alpha$ edges $\{u_{1},x_{1}\},\ldots,\{u_{1},x_{\alpha}\}$, where $\alpha:=s-t-a^{2}-\mathbf{1}_{\{2\nmid n\}}a$.
 Then $e(G_{1})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{1})=s$.
 Secondly, let $G_{2}$ be the graph with the same vertex partition.
-Unlike the previous construction, we now pick $2s$ vertices $x_{1},y_{1},\ldots,x_{s},y_{s}\in A$ and one vertex $u\in B$, add the edges $\{x_{i},y_{i}\}$ into $A$, and delete the edges $\{u,x_{1}\},\ldots,\{u,x_{\alpha}\}$ between $A$ and $B$.
+Unlike the previous construction, we now pick $2s$ vertices $x_{1},y_{1},\ldots,x_{s},y_{s}\in A$ and one vertex $u\in B$, add the $s$ edges $\{x_{i},y_{i}\}$, for $i\in\{1,\ldots,s\}$ into $A$, and delete the $\alpha$ edges $\{u,x_{1}\},\ldots,\{u,x_{\alpha}\}$ between $A$ and $B$.
 Similarly, we have $e(G_{2})=\lfloor n^{2}/4\rfloor+t$ and $\tau_{3}(G_{2})=s$.
@@
 If $\tau_{3}(G)\geq s$ and $n$ is sufficiently large, then $G$ contains at least $\min\{t(G_{1}),t(G_{2})\}$ triangles.
-An independent work was also studied by Liu and Mubayi [32, Theorem 1.4].
-In other words, there are two ways to embed $s$ disjoint edges into the partition in order to minimize the number of triangles, i.e., embedding $s-1$ or $s$ edges into the larger vertex part, and other edges into the smaller part.
-Recently, Li, Feng and Peng [27] presented a spectral version of the result of Xiao and Katona by showing that if $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq 2$, then $G$ contains at least $n-3$ triangles.
-This bound is the best possible and it provides a spectral stability of Ning–Zhai’s result in Theorem 1.3.
-Next, we shall consider the general case $\tau_{3}(G)\geq s$ for every $s\in\mathbb{N}$. The third main result in this paper reveals an asymptotically spectral extension on Theorem 2.4.
+An independent work was also studied by Liu and Mubayi [34, Theorem 1.4].
+In other words, there are two ways to embed $s$ disjoint edges in the partition in order to minimize the number of triangles, that is, embedding s1s-1 or $s$ edges in the larger vertex part, and other edges in the smaller part.
+Recently, Li, Feng and Peng [29] presented a spectral version of the result of Xiao and Katona [59] by showing that if $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq 2$, then $G$ contains at least n3n-3 triangles.
+This bound is the best possible and provides a spectral stability of the Ning–Zhai result in Theorem 1.3.
+Next, we shall consider the general case $\tau_{3}(G)\geq s$ for every $s\in\mathbb{N}$. The third main result in this paper reveals an asymptotically spectral extension of Theorem 2.4.
 Theorem 2.5.
-If $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, then for every $n\geq 113s^{2}$, there are at least $s{n}/{2}-5s^{2}$ triangles in $G$.
+If $G$ is a graph with $n\geq 28s^{2}$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, then there are at least $\frac{1}{2}sn-5s^{2}$ triangles in $G$.
 Remark.
-Theorem 2.5 gives a stability of Theorem 2.3.
-By applying the fact $\lambda(G)\geq 2e(G)/n$, we can see that $e(G)>e(T_{n,2})$ implies $\lambda(G)>\lambda(T_{n,2})$.
-As a by-product of Theorem 2.5, we obtain that if $e(G)>e(T_{n,2})$ and $\tau_{3}(G)\geq s$, then $G$ has $s{n}/{2}-O(1)$ triangles.
-This result could be viewed as an asymptotic version of Theorem 2.4, since $t(G_{1})\approx t(G_{2})\approx s{n}/{2}-O(1)$.
-With additional efforts, one can improve the constant in Theorem 2.5, but it seems difficult to determine the exact bound or the spectral extremal graphs corresponding to the edge-extremal graphs $G_{1}$ and $G_{2}$ in the general case.
+Theorem 2.5 gives the stability of Theorem 2.3.
+Applying the fact $\lambda(G)\geq\frac{2e(G)}{n}$, we see that $e(G)>e(T_{n,2})$ implies $\lambda(G)>\lambda(T_{n,2})$.
+As a by-product of Theorem 2.5, we find that if $e(G)>e(T_{n,2})$ and $\tau_{3}(G)\geq s$, then $G$ contains $\frac{1}{2}sn-O(1)$ triangles.
+This result could be viewed as an asymptotic version of Theorem 2.4, since $t(G_{1})\approx t(G_{2})\approx\frac{1}{2}sn-O(1)$.
+With additional efforts, the constant term of Theorem 2.5 can be slightly improved, but it seems difficult to establish the exact bound and determine the spectral extremal graphs corresponding to the extremal graphs $G_{1}$ and $G_{2}$ in Theorem 2.4.
+2.3 Our approach and organization
 Our approach.
-The main idea in our proofs of Theorems 2.2, 2.3 and 2.5 uses a very simple yet surprisingly powerful technique known as the supersaturation-stability; see Theorem 3.3.
-Informally, every graph with few triangles can be made bipartite by deleting a quite small number of edges.
-The word “quite small” means a small constant number of edges, say $6q$ in Lemma 4.1, which is independent of the order $n$.
-This is slightly different from the classical stability method, which requires us to delete roughly $\varepsilon n^{2}$ edges with a small number $\varepsilon>0$.
-Although this might sound somewhat vague, we will make it more precise in Section 4.
-Unlike the traditional stability method, it is sufficient to utilize the condition $n\geq 300q^{2}$ only, instead of requiring that $q$ is fixed and $n$ is sufficiently large (or a tower-type function on $q$), since our proof avoids the use of the celebrated triangle removal lemma***The lemma says that for each $\varepsilon>0$, there is $\delta>0$ such that every graph on $n$ vertices with at most $\delta n^{3}$ triangles can be made triangle-free by removing at most $\varepsilon n^{2}$ edges.
+The main idea in our proofs of Theorems 2.2, 2.3 and 2.5 uses a simple yet surprisingly powerful technique known as the supersaturation-stability method.
+Informally, every graph with few triangles can be made bipartite by deleting a relatively small number of edges.
+Here, “relatively small” means a constant number of edges independent of $n$; see Theorem 3.6.
+This is different from the classical stability method, which typically requires the removal of roughly $\varepsilon n^{2}$ edges for some small $\varepsilon>0$.
+Although this description may sound somewhat vague, we will make it more precise in Section 4.
+Unlike the traditional stability method, our method requires only the condition $n\geq 121q^{2}$, instead of requiring that $q$ is fixed and $n$ is sufficiently large (or a tower-type function in $q$).
+This is because our proof avoids the use of the celebrated triangle removal lemma***The lemma says that for each $\varepsilon>0$, there is $\delta>0$ such that every graph on $n$ vertices with at most $\delta n^{3}$ triangles can be made triangle-free by removing at most $\varepsilon n^{2}$ edges.
 That is, a graph with a subcubic number of triangles can be made triangle-free by removing a subquadratic number of edges; see [13]..
-This allows us to put $q$ in a wide range that can be growth with $n$.
-Hence, we substantially improve the lower bound on the order $n$ of a graph with respect to $q$.
-We believe that the supersaturation-stability method used in this paper may be interesting in its own right, and perhaps could have other applications elsewhere, just as the classical stability approach has now seen widespread use.
-Last but not least, apart from using the supersaturation-stability, we shall make use of the spectral technique due to Cioabă, Feng, Tait and Zhang [9] as well as the double-eigenvector technique applied by Fang, Tait and Zhai [19].
+Consequently, $q$ is allowed to vary over a wide range and may even grow with $n$.
+Hence, we significantly improve the lower bound on the order $n$ of a graph with respect to $q$.
+We believe that the supersaturation-stability method used in this paper may be of independent interest and could potentially find further applications elsewhere, much like the classical stability approach, which has now seen widespread use.
+Last but not least, in addition to the supersaturation-stability method, we also employ the spectral technique developed by Cioabă, Feng, Tait and Zhang [9], as well as the double-eigenvector technique applied by Fang, Tait and Zhai [19].
 Organization.
-In Section 3, we shall introduce some preliminaries, including the spectral triangle counting lemma and the supersaturation-stability; see Lemma 3.2 and Theorem 3.3, respectively.
-Intuitively it is clear that if a graph $G$ has large spectral radius and it is far from being bipartite, then it is unlikely that all the triangles of $G$ are located on few edges of $G$.
-To warm up, we shall give a “strong” stability result for graphs with few triangles and large spectral radius; see Theorem 3.4.
-In Section 4, we will outline the sketch of our method and then we present the detailed proof of Theorem 2.2.
-In Section 5, we shall use the double-eigenvector technique to give the proof of Theorem 2.3.
-The proof of Theorem 2.5 will be provided in Section 6.
-At the end of this paper, we will propose some related spectral extremal graph problems for interested readers.
+In Section 3, we show the tightness of $q$ in Theorem 2.2 and introduce preliminaries for counting triangles.
+As a warm up, we give a spectral stability result for graphs with few triangles; see Theorem 3.6.
+In Section 4, we outline the main ideas of our method and then present the proof of Theorem 2.2.
+In Section 5, we employ the double-eigenvector technique to prove Theorem 2.3.
+The proof of Theorem 2.5 is provided in Section 6.
+Finally, we propose several related spectral extremal problems for readers.
 Notation.
-We usually write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where each $e_{i}$ is a two-element subset of $V$.
-Here, we admit $|V|=n$ and $|E|=m$.
+We write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where we admit $|V|=n$ and $|E|=m$.
 If $S\subseteq V$ is a subset of the vertex set, then $G[S]$ denotes the subgraph of $G$ induced by $S$, i.e., the graph on $S$ whose edges are those edges of $G$ with both endpoints in $S$.
 For abbreviation, we denote $e(S)=e(G[S])$.
-We will write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and similarly, we write $e(S,T)$ for the number of edges of $G[S,T]$.
-For every $s,t\in\mathbb{N}$, we write $K_{s,t}$ for the complete bipartite graph with parts of sizes $s$ and $t$.
+We write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and we write $e(S,T)=e(G[S,T])$.
+We write $K_{s,t}$ for the complete bipartite graph with parts of sizes $s$ and $t$.
 Let $N(v)$ be the set of vertices adjacent to a vertex $v$ and let $d(v)=|N(v)|$.
-Moreover, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$ for simplicity.
-We will write $\lambda(G)$ for the spectral radius of $G$ and $t(G)$ for the number of triangles in $G$.
+For simplicity, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$.
+We write $\lambda(G)$ for the spectral radius of $G$ and $t(G)$ for the number of triangles in $G$.
 Sometimes, we write $\lambda$ and $t$ for short.
@@
 For a vertex set $S\subseteq V(G)$, we write $G\setminus S$ for the subgraph of $G$ by deleting all vertices of $S$ and their incident edges.
-Let $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})^{\top}$ be a Perron vector of $G$, i.e., a non-negative eigenvector corresponding to $\lambda(G)$.
-By scaling, we may assume further that $\max\{x_{v}:v\in V(G)\}=1$.
-For a subset $S\subseteq V(G)$, we denote $x_{S}=\sum_{v\in S}x_{v}$ for notational convenience.
+Let $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be a Perron vector of $G$, i.e., a non-negative eigenvector corresponding to $\lambda(G)$.
+By scaling, we may further assume that $\max\{x_{v}:v\in V(G)\}=1$.
+For a subset $S\subseteq V(G)$, we denote $x_{S}=\sum_{v\in S}x_{v}$.
 In particular, we have $x_{V}=\sum_{v\in V(G)}x_{v}$.
 3 Preliminaries
-Note that in the case $q=1$, we have $Y_{n,2,1}=T_{n,2,1}$.
-Lemma 3.1 (See [25]).
-(a) If $n$ is even, then $\lambda(Y_{n,2,1})$ is the largest root of
-$f(x)=x^{3}-x^{2}-(n^{2}x)/4+{n^{2}}/{4}-n.$
-(b) If $n$ is odd, then $\lambda(Y_{n,2,1})$ is the largest root of
-$g(x)=x^{3}-x^{2}+{(1-n^{2})x}/{4}+{n^{2}}/{4}-n+{3}/{4}.$
-Consequently, we have
-$\lambda^{2}(Y_{n,2,1})>\left\lfloor{n^{2}}/{4}\right\rfloor+2.$
+Recall that $Y_{n,2,q}$ is a graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the partite set of size $\lceil n/2\rceil$.
+In what follows, we give an estimate on the spectral radius of $Y_{n,2,q}$.
+Lemma 3.1.
+(a) If $n$ is even, then $\lambda(Y_{n,2,q})$ is the largest root of
+$f_{1}(x)=x^{3}-x^{2}-(n^{2}x)/4+n^{2}/4-qn.$
+(b) If $n$ is odd, then $\lambda(Y_{n,2,q})$ is the largest root of
+$f_{2}(x)=x^{3}-x^{2}-(n^{2}x)/4+x/4+n^{2}/4-1/4+q-qn.$
+Consequently, we have $\sqrt{\left\lfloor{n^{2}}/{4}\right\rfloor+2q}<\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ for every $n\geq 4q$.
+Proof.
+(a) Let $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be the Perron eigenvector of $\lambda(Y_{n,2,q})$.
+Assume that $V(Y_{n,2,q})=X_{1}\cup X_{2}\cup Y$, where $X_{1}=\{u_{1},v_{1},u_{2},v_{2},\ldots,u_{q},v_{q}\}$ is the set of vertices of the matching of size $q$, $X_{1}\cup X_{2}$ and $Y$ are partite sets of $K_{\frac{n}{2},\frac{n}{2}}$.
+Since two vertices with the same neighborhood have the same value in corresponding coordinates of $\bm{x}$, we may assume that $x_{u_{i}}=x_{v_{i}}=a$ for $1\leq i\leq q$, $x_{u}=b$ for each $u\in X_{2}$, and $x_{w}=c$ for each $w\in Y$.
+Then
+$\begin{cases}\lambda a=a+\frac{n}{2}c,\\ \lambda b=\frac{n}{2}c,\\ \lambda c=2qa+(\frac{n}{2}-2q)b.\end{cases}$
+Thus, $\lambda(Y_{n,2,q})$ is the largest eigenvalue of
+$B_{1}=\begin{bmatrix}1&0&\frac{n}{2}\\ 0&0&\frac{n}{2}\\ 2q&\frac{n}{2}-2q&0\end{bmatrix}$.
+By calculation, we know that $\lambda(Y_{n,2,q})$ is the largest root of
+$f_{1}(x):=\det(xI-B_{1})=x^{3}-x^{2}-(n^{2}x)/4+n^{2}/4-qn.$
+(b) Let $V(Y_{n,2,q})=X_{1}\cup X_{2}\cup Y$ be a partition, where $X_{1}=\{u_{1},v_{1},\ldots,u_{q},v_{q}\}$ spans a matching with $q$ edges, $X_{1}\cup X_{2}$ and $Y$ are partite sets of $K_{\frac{n+1}{2},\frac{n-1}{2}}$ satisfying $|X_{1}|+|X_{2}|=\frac{n+1}{2}$ and $|Y|=\frac{n-1}{2}$.
+A similar argument yields that $\lambda(Y_{n,2,q})$ is the largest eigenvalue of
+$B_{2}=\begin{bmatrix}1&0&\frac{n-1}{2}\\ 0&0&\frac{n-1}{2}\\ 2q&\frac{n+1}{2}-2q&0\end{bmatrix}$.
+Thus, $\lambda(Y_{n,2,q})$ is the largest root of
+$f_{2}(x):=\det(xI-B_{2})=x^{3}-x^{2}+x/4-(n^{2}x)/4+n^{2}/4-1/4-qn+q.$
+Moreover, it is easy to check that
+$f_{1}(\sqrt{n^{2}/4+2q})=2q\sqrt{n^{2}/4+2q}-qn-2q<0$
+and
+$f_{2}(\sqrt{(n^{2}-1)/4+4})=2q\sqrt{(n^{2}-1)/4+2q}-qn-q<0.$
+Consequently, we get $\sqrt{\lfloor n^{2}/4\rfloor+4}<\lambda(Y_{n,2,q})$.
+Similarly, we can verify that $f_{1}(\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}})>0$ and $f_{2}(\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}})>0$, which implies that $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$, as desired.
+∎
+Similar to the proof of Lemma 3.1, we prove the following proposition.
+Proposition 3.2.
+For any $q\geq 2$, we have $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
+Proof.
+We start the proof in two cases. (a) If $n$ is even, by a simple computation as in the proof of Lemma 3.1, we can obtain that $\lambda(T_{n,2,q})$ is the largest root of
+$h_{1}(x)$ $:=x^{4}-x^{3}-(n^{2}x^{2})/4-qx^{2}+(n^{2}x)/4-(nx)/2-nqx$
+$\quad-(nq)/2+(n^{2}q)/4-(nq^{2})/2.$
+Let $f_{1}(x)$ be the polynomial defined in Lemma 3.1. It follows that
+$xf_{1}(x)-h_{1}(x)=qx^{2}+(nx)/2+(nq^{2})/2+(nq)/2-(n^{2}q)/4.$
+It is easy to see that for every $x>\frac{n}{2}$, we have
+$xf_{1}(x)-h_{1}(x)>\tfrac{n}{2}f_{1}(\tfrac{n}{2})-h_{1}(\tfrac{n}{2})=\tfrac{1}{4}(n^{2}+2q(1+q)n)>0.$
+We conclude that $xf_{1}(x)>h_{1}(x)$ for every $x>\frac{n}{2}$.
+Note that $\lambda(Y_{n,2,q})$ is the largest root of $xf_{1}(x)$.
+Combining with $\lambda(Y_{n,2,q}),\lambda(T_{n,2,q})>\frac{n}{2}$, we obtain $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
+(b) If $n$ is odd, then $\lambda(T_{n,2,q})$ is the largest root of
+$h_{2}(x)$ $:=x^{4}-x^{3}+x^{2}/4-(n^{2}x^{2})/4-qx^{2}+x/4-(nx)/2+(n^{2}x)/4+qx-nqx$
+$\quad+q/4-(nq)/2+(n^{2}q)/4+q^{2}/2-(nq^{2})/2.$
+Let $f_{2}(x)$ be the polynomial defined in Lemma 3.1. Then
+$xf_{2}(x)-h_{2}(x)=qx^{2}+(n-1)x/2+(nq)/2-(n^{2}q)/4+(n-1)q^{2}/2-q/4.$
+For every $x>\frac{n-1}{2}$, it follows that
+$xf_{2}(x)-h_{2}(x)>\tfrac{n-1}{2}f_{2}(\tfrac{n-1}{2})-h_{2}(\tfrac{n-1}{2})=\tfrac{1}{4}(n-1)(2q^{2}+n-1)>0.$
+In this case, observe that $\lambda(Y_{n,2,q}),\lambda(T_{n,2,q})>\frac{n-1}{2}$, so we get $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$. ∎
+3.1 The tightness of $q$ in Theorem 2.2
+In this section, we provide an example showing that the bound of $q$ in Theorem 2.2 is tight up to a constant factor.
+Recall that $Y_{n,2,q}$ and $T_{n,2,q}$ are obtained from $T_{n,2}$ by adding a matching, and a star with $q$ edges into the partite set of size $\lceil n/2\rceil$, respectively.
+Example 3.3.
+If $q=0.8\sqrt{n}$ and $n\geq 110$ are integers, then $\lambda(T_{n,2,q-1})>\lambda(Y_{n,2,q})$. However, the graph $T_{n,2,q-1}$ contains exactly $(q-1)\lfloor n/2\rfloor$ triangles.
+Proof.
+We partition the proof into two cases. (a) If $n$ is even, by a similar argument as in the proof of Lemma 3.1, we know that $\lambda(T_{n,2,q-1})$ is the largest root of
+$g_{1}(x)$ $:=x^{4}-x^{3}-(n^{2}x^{2})/4-qx^{2}+x^{2}+(n^{2}x)/4+(nx)/2-nqx$
+$\quad-(n^{2}/4)+(nq)/2+(n^{2}q)/4-(nq^{2})/2.$
+Let $f_{1}(x)$ be defined in Lemma 3.1. Since $q=0.8\sqrt{n}$, by computation, we see that
+$xf_{1}(x)-g_{1}(x)=(0.8\sqrt{n}-1)x^{2}-0.5nx-0.2n^{5/2}+0.57n^{2}-0.4n^{3/2}.$
+For every $n\geq 4$ and x>n/2x>{n}/{2}, it follows that
+$\frac{\mathrm{d}}{\mathrm{d}x}(xf_{1}(x)-g_{1}(x))=(1.6\sqrt{n}-2)x-0.5n>0.$
+Then for every $x>n/2$, we have
+$xf_{1}(x)-g_{1}(x)\geq\tfrac{n}{2}f_{1}(\tfrac{n}{2})-g_{1}(\tfrac{n}{2})=0.07n^{2}-0.4n^{3/2}>0.$
+We conclude that $xf_{1}(x)>g_{1}(x)$ for every $x>n/2$ and $n\geq 33$.
+Note that $\lambda(Y_{n,2,q})$ and $\lambda(T_{n,2,q-1})$ are the largest roots of $xf_{1}(x)$ and $g_{1}(x)$, respectively.
+Since $\lambda(Y_{n,2,q})>\frac{n}{2}$ and $\lambda(T_{n,2,q-1})>\frac{n}{2}$, we obtain $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q-1})$.
+(b) If $n$ is odd, by computation, we get that $\lambda(T_{n,2,q-1})$ is the largest root of
+$g_{2}(x)$ $=x^{4}-x^{3}+(5x^{2})/4-(n^{2}x^{2})/4-qx^{2}-(3x)/4+(nx)/2+(n^{2}x)/4+qx-nqx$
+$\quad+1/4-n^{2}/4-(3q)/4+(nq)/2+(n^{2}q)/4+q^{2}/2-(nq^{2})/2.$
+Let $f_{2}(x)$ be determined in Lemma 3.1. Substituting $q=0.8\sqrt{n}$, we compute that
+[…]
+[…]
+For every $x>\frac{n-1}{2}$, we have
+$\frac{\mathrm{d}}{\mathrm{d}x}(xf_{2}(x)-g_{2}(x))=(1.6\sqrt{n}-2)x-0.5(n-1)>0.$
+It follows that for every $n\geq 110$ and $x>\frac{n-1}{2}$,
+$xf_{2}(x)-g_{2}(x)>\tfrac{n-1}{2}f_{2}(\tfrac{n-1}{2})-g_{2}(\tfrac{n-1}{2})=0.07n^{2}-0.8n^{3/2}+0.68n+0.8\sqrt{n}-0.75>0.$
+Thus, we get $xf_{2}(x)>g_{2}(x)$ for every $x>\frac{n-1}{2}$.
+It is easy to see that $\lambda(Y_{n,2,q})>\frac{n-1}{2}$ and $\lambda(T_{n,2,q-1})>\frac{n-1}{2}$.
+So we conclude that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q-1})$.
+∎
+3.2 Counting the number of triangles
 In 2007, Bollobás and Nikiforov [7] established some powerful results on counting the number of triangles (cliques) of a graph in terms of the spectral radius.
@@
 The equality holds if and only if $G$ is a complete bipartite graph.
-The inequality can be written as the following versions:
+As a consequence, Lemma 3.4 implies an interesting result of Nosal [50], which states that every $m$-edge graph $G$ with $\lambda(G)>\sqrt{m}$ contains a triangle; see [41, 47] for related results.
+The inequality in Lemma 3.4 can be rephrased as the following versions:
 $\lambda^{3}\leq 3t+m\lambda\quad\Leftrightarrow\quad{t\geq\frac{1}{3}\lambda(\lambda^{2}-m)}\quad\Leftrightarrow\quad m\geq\lambda^{2}-\frac{3t}{\lambda}.$
-This inequality was initially obtained by Bollobás and Nikiforov [7, Theorem 1], and it was independently discovered by Cioabă, Feng, Tait and Zhang [9].
-The case of equality was recently characterized by Ning and Zhai [45].
+This inequality was initially obtained by Bollobás and Nikiforov [7, Theorem 1], and it was independently discovered by Cioabă, Feng, Tait and Zhang [9], as well as Ning and Zhai [48].
+The case of equality was characterized by Ning and Zhai [48].
 Moon–Moser’s inequality states that if $G$ has $n$ vertices and $m$ edges, then
@@
 We refer to [4, p.
-297] and [33, p.
+297] and [36, p.
 443] for detailed proofs.
@@
 Next, we present a generalization for a graph with less than $n^{2}/4$ edges whenever this graph is far from being bipartite.
-More precisely, for $\varepsilon>0$, we say that a graph $G$ is $\varepsilon$-far from being bipartite if $G^{\prime}$ is not bipartite for every subgraph $G^{\prime}\subseteq G$ with $e(G^{\prime})>e(G)-\varepsilon$.
-In other words, if $G$ is $\varepsilon$-far from being bipartite, then we need to delete at least $\varepsilon$ edges from $G$ to make it bipartite.
+More precisely, for $\varepsilon>0$, we say that a graph $G$ is $\varepsilon$-far from being bipartite if any subgraph $G^{\prime}$ of $G$ with more than $e(G)-\varepsilon$ edges is not bipartite.
+In other words, if $G$ is $\varepsilon$-far from being bipartite, then we must delete at least $\varepsilon$ edges from $G$ to make it bipartite.
 Theorem 3.5 (See [2]).
@@
 $t(G)\geq\frac{n}{6}\left(m+\varepsilon-\frac{n^{2}}{4}\right).$
+For the convenience of the reader, we include a proof here.
+Proof.
+For each vertex $v\in V(G)$, let $N_{v}=N(v)$ and $N_{v}^{c}=V(G)\setminus N(v)$.
+Since $G$ is $t$-far from being bipartite, we must have $e(N_{v})+e(N_{v}^{c})\geq\varepsilon$ for every $v\in V(G)$.
+Indeed, if $e(N_{v})+e(N_{v}^{c})<\varepsilon$ for some $v$, then deleting all edges inside $N_{v}$ and $N_{v}^{c}$ yields a bipartite subgraph with more than $e(G)-\varepsilon$ edges, a contradiction.
+For each $v\in V(G)$, we have
+$\sum_{w\in N_{v}^{c}}d(w)=2e(N_{v}^{c})+e(N_{v}^{c},N_{v})=e(N_{v}^{c})+m-e(N_{v})\geq m+\varepsilon-2e(N_{v}).$
+Summing this inequality over all vertices $v\in V(G)$ gives
+$\sum_{v\in V}\sum_{w\in N_{v}^{c}}d(w)\geq mn+\varepsilon n-2\sum_{v\in V}e(N_{v})=mn+\varepsilon n-6t(G),$
+where we used the identity $3t(G)=\sum_{v\in V}e(N_{v})$. On the other hand, we get
+$\sum_{v\in V}\sum_{w\in N_{v}^{c}}d(w)=\sum_{v\in V}\left(2m-\sum_{w\in N_{v}}d(w)\right)=2mn-\sum_{w\in V}d^{2}(w).$
+Combining the two inequalities, we obtain
+$6t(G)\geq\varepsilon n-mn+\sum_{w\in V}d^{2}(w)\geq\varepsilon n-mn+\frac{4m^{2}}{n},$
+where the last inequality follows by Cauchy–Schwarz’s inequality. Note that ${4m^{2}}/{n}\geq 2mn-{n^{3}}/{4}$. Substituting this into the previous inequality gives the desired bound. ∎
 The supersaturation on triangles is usually deduced from Moon–Moser’s inequality.
@@
 Thus, if $G$ has at least ${n^{2}}/{4}+q$ edges, then it is $q$-far from being bipartite.
-Theorem 3.3 implies that $G$ contains at least ${qn}/{3}$ triangles.
+Theorem 3.5 implies that $G$ contains at least qn/3{qn}/{3} triangles.
 In particular, setting $q=\varepsilon n^{2}$, we obtain that $G$ has at least ${\varepsilon n^{3}}/{3}$ triangles.
 For dealing with some extremal graph problems, it is a great advantage that applying Theorem 3.5 can help us to get rid of the use of triangle removal lemma so that we do not require the order of a graph to be sufficiently large.
-This approach could be considered to be a key innovation in our paper, and it was proven to be effective and feasible in many extremal graph problems.
-We refer to [26, 27] for more related applications.
-In what follows, we provide a warm-up practice that has its independent interest.
-We concentrate on explaining the main ideas which we believe have wide applicability.
+This approach could be considered to be a key ingredient in our paper, and it was proven to be effective and feasible in many extremal graph problems.
+For other related applications, we refer the reader to [28, 29].
+3.3 A spectral stability for graphs with few triangles
+In what follows, we provide a warm-up practice that has its independent interest. We concentrate on explaining the main ideas which we believe have wide applicability.
 Theorem 3.6.
 If $G$ is a graph on $n$ vertices with $\lambda(G)\geq n/2$ and $t(G)\leq kn/2$, then
-$e(G)\geq\frac{n^{2}}{4}-3k,$
+$e(G)>\frac{n^{2}}{4}-3k,$
 and there exists a vertex partition $V(G)=S\cup T$ such that
@@
 Proof.
-First of all, if $G$ is a complete bipartite graph, then $e(G)\leq\lfloor n^{2}/4\rfloor$.
-Moreover, we have $e(G)=\lambda^{2}\geq n^{2}/4$.
-It follows that $n$ is even and $G=K_{\frac{n}{2},\frac{n}{2}}$.
+First, if $G$ is a complete bipartite graph, then $e(G)\leq\lfloor n^{2}/4\rfloor$.
+Since $e(G)=\lambda^{2}(G)\geq n^{2}/4$, it follows that $n$ is even and $G=K_{\frac{n}{2},\frac{n}{2}}$.
 In this case, the desired result holds immediately.
-Now, we consider the case that $G$ is not a complete bipartite graph, then the inequality in Lemma 3.2 holds strictly.
+Now suppose that $G$ is not a complete bipartite graph.
+Then the inequality in Lemma 3.4 is strict.
 So we get $e(G)>\lambda^{2}-(3t)/\lambda\geq\lambda^{2}-(6t)/n\geq n^{2}/4-3k$.
 Theorem 3.5 implies that $G$ is not $6k$-far from being bipartite.
-Otherwise, if $G$ is $6k$-far from being bipartite, then $t(G)\geq{n}/{6}\cdot(e(G)+6k-n^{2}/4)>kn/2$, which contradicts with the assumption.
+Otherwise, if $G$ is $6k$-far from being bipartite, then $t(G)\geq{n}/{6}\cdot(e(G)+6k-n^{2}/4)>kn/2$, which contradicts the assumption.
 Hence $G$ is not $6k$-far from being bipartite.
-Then there exists a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6k$.
-Moreover, we can get $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4-9k$.
-Consequently, we obtain that $n/2-3\sqrt{k}<|S|,|T|<n/2+3\sqrt{k}$.
+Consequently, there exists a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6k$.
+Moreover, we get $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4-9k$.
+It follows that $n/2-3\sqrt{k}<|S|,|T|<n/2+3\sqrt{k}$.
 Note that $e(K_{|S|,|T|})-e(S,T)<|S||T|-(n^{2}/4-9k)\leq 9k$.
-In other words, there are less than $9k$ missing edges between $S$ and $T$.
+In other words, there are fewer than $9k$ missing edges between $S$ and $T$.
 Consequently, we get $\delta(G)>(n/2-3\sqrt{k})-9k\geq n/2-12k$.
-Note that $e(S)+e(T)<6k$.
-Then it follows that $\Delta(G)<(n/2+3\sqrt{k})+6k\leq n/2+9k$, as desired.
+Finally, since $e(S)+e(T)<6k$, we have $\Delta(G)<(n/2+3\sqrt{k})+6k\leq n/2+9k$, as desired.
 ∎
 4 Proof of Theorem 2.2
-Note that the case $q=1$ in Problem 2.1 was proved by Li, Lu and Peng [25].
-Next, we shall consider the case $q\geq 2$ and $n\geq 300q^{2}$ although our proof works for $q=1$.
-To avoid unnecessary computations, we did not make much efforts to optimize the bound on the order of a graph.
+The case $q=1$ was proved by Li, Lu and Peng [27].
+Next, we consider the case $2\leq q\leq\frac{1}{11}\sqrt{n}$ although our proof works for $q=1$.
 Throughout this section, let $G$ be a graph on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$.
-Suppose on the contrary that $G$ contains less than $q\lfloor{n}/{2}\rfloor$ triangles.
+Suppose on the contrary that $t(G)<q\lfloor{n}/{2}\rfloor$.
 In the sequel, we will deduce a contradiction.
-The essential observation in our proof attributes to the supersaturation-stability as described in Theorem 3.3.
-To start the proof of Theorem 2.2, we outline the main steps as follows.
-  • Firstly, we establish a bipartition of vertices of $G$.
-Using Lemma 3.2, we get that $G$ has more than $n^{2}/4-3q$ edges.
-The supersaturation-stability in Theorem 3.3 implies that one can delete $6q$ edges from $G$ to make it bipartite.
-So we get a partition $V(G)=S\cup T$ with $e(S)+e(T)<6q$ and $|S|,|T|\approx n/2\pm 3\sqrt{q}$.
-  • Secondly, we shall show further that $e(S)+e(T)\leq q$.
-Moreover, the minimum degree of $G$ is at least $n/2-4q$; see Lemmas 4.2 and 4.3.
-Let $\mathbf{x}\in\mathbb{R}^{n}$ be the Perron vector of $G$ satisfying $\max\{x_{v}:v\in V(G)\}=1$.
-Then we will prove in Lemma 4.4 that ${x}_{u}\geq 1-{30q}/{n}$ for every $u\in V(G)$.
-  • Finally, we prove in Lemma 4.6 that the bipartition of $G$ is balanced, i.e., $\bigl{|}|S|-|T|\bigr{|}\leq 1$.
-Using the above lemmas, we shall prove that $G$ has at least $q\lfloor n/2\rfloor$ triangles, or $G$ satisfies $\lambda(G)<\lambda(Y_{n,2,q})$.
-This leads to a contridiction.
+The key insight of our proof lies in the application of the supersaturation-stability (Theorem 3.5).
+We outline the main steps as follows.
+  • Using Lemma 3.4, we get that $G$ contains more than $\lfloor n^{2}/4\rfloor-q$ edges.
+Theorem 3.5 implies that we can delete less than $4q$ edges from $G$ to make it bipartite.
+So we get a bipartition $V(G)=S\cup T$ with $e(S)+e(T)<4q$ and $|S|,|T|\approx\frac{n}{2}\pm\sqrt{5q}$.
+  • Secondly, we show that $e(S)+e(T)\leq q$ and $\delta(G)\geq\frac{n}{2}-2q$; see Lemmas 4.2 and 4.3.
+Let $\bm{x}\in\mathbb{R}^{n}$ be the Perron vector of $G$ satisfying $\max\{x_{v}:v\in V(G)\}=1$.
+Then we will prove in Lemma 4.4 that ${x}_{u}\geq 1-{20q}/{n}$ for every $u\in V(G)$.
+  • We prove in Lemma 4.6 that the bipartition of $G$ is balanced, i.e., $\bigl||S|-|T|\bigr|\leq 1$.
+Then we shall prove that $t(G)\geq q\lfloor n/2\rfloor$, or $G$ satisfies $\lambda(G)<\lambda(Y_{n,2,q})$, a contridiction.
 In what follows, we divide the proof into a sequence of lemmas.
@@
 There is a partition $V(G)=S\cup T$ such that
-$e(S)+e(T)<6q$
+$e(S)+e(T)<4q$
 and
-$e(S,T)>\frac{n^{2}}{4}-9q.$
+$e(S,T)>\frac{n^{2}}{4}-5q.$
 Consequently, we get
-$\frac{n}{2}-3\sqrt{q}<|S|,|T|<\frac{n}{2}+3\sqrt{q}.$
+$\frac{n}{2}-\sqrt{5q}<|S|,|T|<\frac{n}{2}+\sqrt{5q}.$
 Proof.
-First of all, we know from Lemma 3.1 that $\lambda^{2}(Y_{n,2,1})>\lfloor n^{2}/4\rfloor+2$.
-Then $\lambda^{2}(G)\geq\lambda^{2}(Y_{n,2,q})>\lfloor n^{2}/4\rfloor+2$.
-Since $t(G)<q\lfloor{n}/{2}\rfloor$, by Lemma 3.2, we have
-$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n}>\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q.$ (1)
-We claim that $G$ is not $6q$-far from being bipartite. Otherwise, if $G$ is $6q$-far from being bipartite, then using Theorem 3.3, we get
-$t(G)\geq\frac{n}{6}\left(e(G)+6q-\frac{n^{2}}{4}\right)>q\left\lfloor\frac{n}{2}\right\rfloor,$
+By Lemma 3.1, we know that $\lambda^{2}(G)\geq\lambda^{2}(Y_{n,2,q})>\lfloor n^{2}/4\rfloor+2q$. Since $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)>\frac{n}{2}$, by Lemma 3.4, we have
+$e(G)\geq\lambda^{2}(G)-\frac{3t(G)}{\lambda(G)}>\lambda^{2}(G)-\frac{6t(G)}{n}>\left\lfloor\frac{n^{2}}{4}\right\rfloor-q.$ (1)
+We claim that $G$ is not $4q$-far from being bipartite. Otherwise, if $G$ is $4q$-far from being bipartite, then using Theorem 3.5, we get
+$t(G)\geq\frac{n}{6}\left(e(G)+4q-\frac{n^{2}}{4}\right)>q\left\lfloor\frac{n}{2}\right\rfloor,$
 which contradicts with the assumption.
-So $G$ is not $6q$-far from being bipartite.
-That is to say, we can delete less than $6q$ edges from $G$ to make it bipartite, i.e., there exists a vertex partition $V(G)=S\cup T$ such that
-$e(S)+e(T)<6q.$
+So $G$ is not $4q$-far from being bipartite.
+That is to say, we can delete less than $4q$ edges from $G$ to make it bipartite, i.e., there exists a vertex partition $V(G)=S\cup T$ such that
+$e(S)+e(T)<4q.$
 From the above discussion, we get
-$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-9q.$
-Without loss of generality, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq{n}/{2}-3\sqrt{q}$. Then $|T|=n-|S|\geq{n}/{2}+3\sqrt{q}$. It yields that
-$e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-3\sqrt{q}\right)\left(\frac{n}{2}+3\sqrt{q}\right)=\frac{n^{2}}{4}-9q,$
-a contradiction. Thus, we must have ${n}/{2}-3\sqrt{q}<|S|,|T|<{n}/{2}+3\sqrt{q}$. ∎
+$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-5q.$
+Without loss of generality, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq\frac{n}{2}-\sqrt{5q}$. Then $|T|=n-|S|\geq\frac{n}{2}+\sqrt{5q}$. It yields that
+$e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-\sqrt{5q}\right)\left(\frac{n}{2}+\sqrt{5q}\right)=\frac{n^{2}}{4}-5q,$
+a contradiction. Thus, we must have $\frac{n}{2}-\sqrt{5q}<|S|,|T|<\frac{n}{2}+\sqrt{5q}$. ∎
 Lemma 4.2.
@@
 Proof.
-Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$.
-Suppose on the contrary that $e(S)+e(T)\geq q+1$.
-We write $K_{|S|,|T|}$ for the complete bipartite graph on the vertex parts $S$ and $T$.
-Note that each edge of $G[S]$ or $G[T]$ can yield at least $\min\{|S|,|T|\}$ triangles in $K_{|S|,|T|}$.
-Recall in Lemma 4.1 that $e(S)+e(T)<6q$.
-So each missing edge between $S$ and $T$ destroys at most $6q$ triangles.
-By Lemma 4.1 again, we have $\min\{|S|,|T|\}>{n}/{2}-3\sqrt{q}$, and $G[S,T]$ misses at most $9q$ edges from the graph $K_{|S|,|T|}$.
-Hence, we conclude
-$t(G)\geq(q+1)\left(\frac{n}{2}-3\sqrt{q}\right)-9q\cdot 6q>(q+1)\frac{n}{2}-58q^{2}\geq q\left\lfloor\frac{n}{2}\right\rfloor,$
+Suppose on the contrary that $e(S)+e(T)=k$, where $k$ is an integer with $q+1\leq k<4q$ by Lemma 4.1.
+For simplicity, we write $K_{S,T}$ for the complete bipartite graph on the vertex parts $S$ and $T$.
+Note that every triangle of $G$ must contain an edge from $G[S]\cup G[T]$, and each edge of $G[S]\cup G[T]$ yields at least $\min\{|S|,|T|\}$ triangles.
+Each missing edge between $S$ and $T$ destroys at most $k$ triangles.
+By Lemma 4.1, we have $\min\{|S|,|T|\}>\frac{n}{2}-\sqrt{5q}$, and $G[S,T]$ misses at most $5q$ edges from $K_{S,T}$.
+We conclude
+$t(G)>k\left(\frac{n}{2}-\sqrt{5q}\right)-5q\cdot k\geq(q+1)\left(\frac{n}{2}-\sqrt{5q}-5q\right)\geq q\left\lfloor\frac{n}{2}\right\rfloor,$
 where the last inequality holds for every $n\geq 20q^{2}$. This contradicts with the assumption $t(G)<q\lfloor n/2\rfloor$. Thus, we have $e(S)+e(T)\leq q$. ∎
 Lemma 4.3.
-We have
-$e(S,T)>\frac{n^{2}}{4}+2-4q$
+With the previous notation, we have
+$e(S,T)>\frac{n^{2}}{4}-2q$
 and
-$\frac{n}{2}-2\sqrt{q}<|S|,|T|<\frac{n}{2}+2\sqrt{q}.$
+$\frac{n}{2}-\sqrt{2q}<|S|,|T|<\frac{n}{2}+\sqrt{2q}.$
 Furthermore, it follows that
-$\frac{n}{2}-4q\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+q+2\sqrt{q}.$
+$\frac{n}{2}-2q\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+q+\sqrt{2q}.$
 Proof.
-By Lemma 4.2, we obtain $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4+2-4q$. By the AM-GM inequality, we can get ${n}/{2}-2\sqrt{q}\leq|S|,|T|\leq{n}/{2}+2\sqrt{q}$.
-By Rayleigh’s formula, it is well-known that $\delta(G)\leq\lambda(G)\leq\Delta(G)$.
+Combining (1) with Lemma 4.2, we obtain $e(S,T)=e(G)-e(S)-e(T)>\lfloor n^{2}/4\rfloor-2q$. By the AM-GM inequality, we can get $\frac{n}{2}-\sqrt{2q}\leq|S|,|T|\leq\frac{n}{2}+\sqrt{2q}$.
+By the Rayleigh formula, it is well-known that $\delta(G)\leq\lambda(G)\leq\Delta(G)$.
 Since $e(S)+e(T)\leq q$, each vertex of $S$ has degree at most $q$ in $G[S]$, and this also holds for the vertices of $T$.
-Then it follows that $\Delta(G)\leq\max\{|S|,|T|\}+q\leq{n}/{2}+2\sqrt{q}+q$.
-Next, we show that $\delta(G)\geq{n}/{2}-4q$.
-Otherwise, if there exists a vertex $v\in V(G)$ with the minimum degree and $d(v)\leq{n}/{2}-4q-{1}/{2}$, then
-$e(G\setminus\{v\})=e(G)-d(v)$$\geq$ $\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q-\left(\frac{n}{2}-4q-\frac{1}{2}\right)$
-$\geq$ $\frac{(n-1)^{2}}{4}+q+2.$
-By Theorem 1.2, we can find $(q+2)\lfloor\frac{n-1}{2}\rfloor$ triangles in $G\setminus\{v\}$.
-This is a contradiction with the assumption $t(G)<q\lfloor{n}/{2}\rfloor$.
-Thus, we must have $\delta(G)\geq{n}/{2}-4q$.
+Then it follows that $\Delta(G)\leq\max\{|S|,|T|\}+q\leq\frac{n}{2}+\sqrt{2q}+q$.
+Next, we show that $\delta(G)\geq\frac{n}{2}-2q$.
+Otherwise, if there exists a vertex $v\in V(G)$ with degree and $d(v)\leq\frac{n}{2}-2q-\frac{1}{2}$, then using (1) gives
+$e(G\setminus\{v\})=e(G)-d(v)$ $\geq$ $\left(\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-q+1\right)-\left(\frac{n}{2}-2q-\frac{1}{2}\right)$
+$\geq$ $\frac{(n-1)^{2}}{4}+q+1.$
+By Theorem 1.2, we can find at least $(q+1)\lfloor\frac{n-1}{2}\rfloor$ triangles in $G\setminus\{v\}$.
+This is a contradiction with the assumption $t(G)<q\lfloor\frac{n}{2}\rfloor$.
+Thus, we must have $\delta(G)\geq\frac{n}{2}-2q$.
 ∎
-Recall that $\mathbf{x}\in\mathbb{R}^{n}$ is the Perron vector of $G$ with $\max\{x_{v}:v\in V(G)\}=1$.
+Recall that $\bm{x}\in\mathbb{R}^{n}$ is the Perron vector of $G$ with $\max\{x_{v}:v\in V(G)\}=1$.
 Lemma 4.4.
 For every $u\in V(G)$, we have
-${x}_{u}>1-\frac{30q}{n}.$
+${x}_{u}>1-\frac{20q}{n}.$
 Proof.
-Let $z\in V(G)$ be a vertex such that ${x}_{z}$ attains the maximum eigenvector entry of $\mathbf{x}$ and ${x}_{z}=1$.
+Let $z\in V(G)$ be a vertex such that ${x}_{z}$ attains the maximum eigenvector entry of $\bm{x}$ and ${x}_{z}=1$.
 Without loss of generality, we may assume that $z\in S$.
 Our proof of this lemma falls naturally into two steps.
-Step 1. For every $u\in S$, we have ${x}_{u}>1-\frac{15q}{n}$.
-Since $e(S)\leq q$ by Lemma 4.2, we get $d_{S}(u)\leq q$. By Lemma 4.3, we have $d_{T}(u)=d(u)-d_{S}(u)\geq({n}/{2}-4q)-q={n}/{2}-5q$. By Lemma 4.3 again, we get $|T|\leq{n}/{2}+2\sqrt{q}$. Then
+Step 1. For every $u\in S$, we have ${x}_{u}>1-\frac{11q}{n}$.
+Using Lemma 4.2 gives $e(S)\leq q$ and $d_{S}(u)\leq q$.
+By Lemma 4.3, we have $d_{T}(u)=d(u)-d_{S}(u)\geq(\frac{n}{2}-2q)-q=\frac{n}{2}-3q$.
+By Lemma 4.3 again, we get $|T|\leq\frac{n}{2}+\sqrt{2q}$.
+Then
 $|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|$ $=$ $-|N_{T}(u)|+|N_{T}(u)\cup N_{T}(z)|$
-$\leq$ $-\left(\frac{n}{2}-5q\right)+\left(\frac{n}{2}+2\sqrt{q}\right)$
-$=$ $5q+2\sqrt{q}.$
+$\leq$ $-\left(\frac{n}{2}-3q\right)+\left(\frac{n}{2}+\sqrt{2q}\right)$
+$=$ $3q+\sqrt{2q}.$
 Hence, we have
-$\lambda{x}_{u}-\lambda{x}_{z}=\sum_{v\sim u}{x}_{v}-\sum_{v\sim z}{x}_{v}\geq-\sum_{v\sim z,v\in T,v\not\sim u}{x}_{v}-\sum_{v\sim z,v\in S}{x}_{v}.$
+$\lambda{x}_{u}-\lambda{x}_{z}=\sum_{v\sim u}{x}_{v}-\sum_{v\sim z}{x}_{v}\geq-\sum_{v\in T,v\sim z,v\not\sim u}{x}_{v}-\sum_{v\in S,v\sim z}{x}_{v}.$
 Applying the fact ${x}_{v}\leq 1$, it follows that
-$\lambda{x}_{u}-\lambda{x}_{z}\geq-\Bigl{(}|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|\Bigr{)}-|N_{S}(z)|\geq-6q-2\sqrt{q}.$
+$\lambda{x}_{u}-\lambda{x}_{z}\geq-\Bigl(|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|\Bigr)-|N_{S}(z)|\geq-4q-\sqrt{2q}.$
 Recall that ${x}_{z}=1$ and $\lambda>{n}/{2}$. Then for each $u\in S$, we have
-${x}_{u}\geq 1-\frac{6q+2\sqrt{q}}{\lambda}>1-\frac{12q+4\sqrt{q}}{n}>1-\frac{15q}{n}.$
-Step 2. For every $u\in T$, we have ${x}_{u}>1-\frac{30q}{n}$.
+${x}_{u}\geq 1-\frac{4q+\sqrt{2q}}{\lambda}>1-\frac{8q+2\sqrt{2q}}{n}>1-\frac{11q}{n}.$
+Step 2. For every $u\in T$, we have ${x}_{u}>1-\frac{20q}{n}$.
 By the previous step, we get
-$\lambda{x}_{u}=\sum_{v\in N(u)}{x}_{v}\geq\sum_{v\in N_{S}(u)}{x}_{v}\geq\left(1-\frac{12q+4\sqrt{q}}{n}\right)d_{S}(u).$
-By Lemma 4.3, we can see that $d(u)\geq\delta(G)\geq{n}/{2}-4q$, which together with $d_{T}(u)\leq q$ gives $d_{S}(u)=d(u)-d_{T}(u)\geq{n}/{2}-5q$. Hence, for $q\geq 2$, we have
-${x}_{u}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})d_{S}(u)}{\lambda}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})(\frac{n}{2}-5q)}{\frac{n}{2}+q+2\sqrt{q}}>1-\frac{30q}{n}.$
+$\lambda{x}_{u}=\sum_{v\in N(u)}{x}_{v}\geq\sum_{v\in N_{S}(u)}{x}_{v}>\left(1-\frac{8q+2\sqrt{2q}}{n}\right)d_{S}(u).$
+By Lemma 4.3, we can see that $d(u)\geq\delta(G)\geq\frac{n}{2}-2q$, which together with $d_{T}(u)\leq q$ gives $d_{S}(u)=d(u)-d_{T}(u)\geq\frac{n}{2}-3q$. Hence, for $q\geq 2$, we have
+${x}_{u}>\frac{(1-\frac{8q+2\sqrt{2q}}{n})d_{S}(u)}{\lambda}\geq\frac{(1-\frac{8q+2\sqrt{2q}}{n})(\frac{n}{2}-3q)}{\frac{n}{2}+q+\sqrt{2q}}\geq 1-\frac{20q}{n}.$
 Combining the above two claims, the desired result follows immediately. ∎
-Using Lemma 4.4, namely, ${x}_{u}=1-O({1}/{n})$, we will show that the bipartition $V(G)=S\cup T$ is balanced (see the forthcoming Lemma 4.6).
-To start with, we fix some notation for convenience.
-Let $K_{|S|,|T|}$ be the complete bipartite graph with partite sets $S$ and $T$.
-Let $G_{1}=G[S]\cup G[T]$ be the graph whose edges consist of the class-edges, and let $G_{2}$ be the graph on $V(G)$ with the missing edges between $S$ and $T$, that is, $E(G_{2})=E(K_{|S|,|T|})\setminus E(G)$.
-Note that $e(G)=e(G_{1})+e(K_{|S|,|T|})-e(G_{2})$.
+Lemma 4.4 gives ${x}_{u}>1-\frac{20q}{n}$, we will show that the bipartition $V(G)=S\cup T$ is balanced (see Lemma 4.6).
+To start with, we fix some notation.
+Let $K_{S,T}$ be the complete bipartite graph with partite sets $S$ and $T$.
+Let $G_{1}=G[S]\cup G[T]$ be the graph whose edges consist of the class-edges, and let $G_{2}$ be the graph on $V(G)$ with the missing edges between $S$ and $T$, that is, $E(G_{2})=E(K_{S,T})\setminus E(G)$.
+Note that $e(G)=e(G_{1})+e(K_{S,T})-e(G_{2})$.
 Lemma 4.5.
 We have
-$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}\leq\frac{120q^{2}}{n(n-60q)}.$
+$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{80q^{2}}{n(n-40q)}.$
 Proof.
 By Lemma 4.2, we get $e(G_{1})\leq q$. By Lemma 4.4, we have
-$\mathbf{x}^{\top}\mathbf{x}>n\left(1-\frac{30q}{n}\right)^{2}>n\left(1-\frac{60q}{n}\right)=n-60q.$
+$\bm{x}^{\mathrm{T}}\bm{x}>n\left(1-\frac{20q}{n}\right)^{2}>n\left(1-\frac{40q}{n}\right)=n-40q.$
 Using Rayleigh’s formula, it follows that
-$\lambda(G)$ $=$ $\frac{\mathbf{x}^{\top}\bigl{(}A(K_{|S|,|T|})+A(G_{1})-A(G_{2})\bigr{)}\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
-$=$ $\frac{\mathbf{x}^{\top}A(K_{|S|,|T|})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}+\frac{\mathbf{x}^{\top}A(G_{1})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
-$<$ $\sqrt{|S||T|}+\frac{2q}{n-60q}.$
+$\lambda(G)$ $=$ $\frac{\bm{x}^{\mathrm{T}}\bigl(A(K_{S,T})+A(G_{1})-A(G_{2})\bigr)\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$
+$=$ $\frac{\bm{x}^{\mathrm{T}}A(K_{S,T})\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}+\frac{\bm{x}^{\mathrm{T}}A(G_{1})\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$
+$<$ $\sqrt{|S||T|}+\frac{2q}{n-40q}.$
 On the other hand, it is familiar that
@@
 Then we have
-$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<2q\left(\frac{1}{n-60q}-\frac{1}{n}\right)=\frac{120q^{2}}{n(n-60q)}.$
+$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<2q\left(\frac{1}{n-40q}-\frac{1}{n}\right)=\frac{80q^{2}}{n(n-40q)}.$
 This completes the proof. ∎
@@
 By Lemma 4.5, we get
-$\frac{1}{n}<\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{120q^{2}}{n(n-60q)}.$
+$\frac{1}{n}<\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{80q^{2}}{n(n-40q)}.$
 This leads to a contradiction for $q\geq 2$ and $n\geq 100q^{2}$. ∎
@@
 By Lemma 4.6, we know that $\bigl||S|-|T|\bigr|\leq 1$.
-Recall that $Y_{n,2,q}$ is the graph on the vertex sets $S$ and $T$ with all edges between $S$ and $T$, together with $q$ pairwise disjoint edges embedded into the vertex part of size $\lceil{n}/{2}\rceil$.
-Clearly, we have $e(Y_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ edges.
+Recall that $Y_{n,2,q}$ is the graph on the vertex sets $S$ and $T$ with all edges between $S$ and $T$, together with $q$ pairwise disjoint edges embedded into the larger partite set.
+Clearly, we have $e(Y_{n,2,q})=\lfloor n^{2}/4\rfloor+q$.
 We denote
@@
 Furthermore, we know from Lemma 4.2 that $|E_{-}|\leq e(S)+e(T)\leq q$.
-By Lemma 4.2 again, we have that $|E_{+}|\leq\lfloor{n^{2}}/{4}\rfloor-e(S,T)+q\leq 5q$.
-By Lemma 4.4, we know that ${x}_{u}\geq 1-{(30q)}/{n}$ for every $u\in V(G)$.
+By Lemma 4.3, we have that $|E_{+}|\leq\lfloor{n^{2}}/{4}\rfloor-e(S,T)+q\leq 3q$.
+By Lemma 4.4, we know that ${x}_{u}>1-{(20q)}/{n}$ for every $u\in V(G)$.
 Therefore, we obtain
-$\lambda(Y_{n,2,q})$$\geq$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(Y_{n,2,q})}{x}_{i}{x}_{j}$
-$=$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(G)}{x}_{i}{x}_{j}+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{+}}{x}_{i}{x}_{j}-\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{-}}{x}_{i}{x}_{j}$
-[…]
-$=$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(|E_{+}|-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|-|E_{-}|\right)$
-$\geq$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(1-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|\right)$
+$\lambda(Y_{n,2,q})$ $\geq$ $\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E(Y_{n,2,q})}{x}_{i}{x}_{j}$
+$=$ $\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E(G)}{x}_{i}{x}_{j}+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E_{+}}{x}_{i}{x}_{j}-\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E_{-}}{x}_{i}{x}_{j}$
+$>$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(|E_{+}|\Bigl(1-\frac{20q}{n}\Bigr)^{2}-|E_{-}|\right)$
+$=$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(|E_{+}|-\frac{40q}{n}|E_{+}|+\frac{(20q)^{2}}{n^{2}}|E_{+}|-|E_{-}|\right)$
+$\geq$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(1-\frac{40q}{n}|E_{+}|+\frac{(20q)^{2}}{n^{2}}|E_{+}|\right)$
 $>$ $\lambda(G),$
-where the last inequality holds for $n\geq 300q^{2}$ since $|E_{+}|\leq 5q$. Therefore, we have $\lambda(Y_{n,2,q})>\lambda(G)$, as desired. ∎
-From the above proof of Lemma 4.7, we can obtain the following proposition.
+where the last inequality holds since $n\geq 120q^{2}$ and $|E_{+}|\leq 3q$. ∎
+From the above proof of Lemma 4.7, we can immediately obtain the following proposition, which can be viewed as a supplement of Example 3.3.
 Proposition 4.8.
-If $n\geq 300q^{2}$ and $G$ is a graph obtained from $T_{n,2}$ by adding $q-1$ edges into the color classes arbitrarily, then $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)<\lambda(Y_{n,2,q})$.
+If $q\leq\frac{1}{11}\sqrt{n}$ and $G$ is a graph obtained from $T_{n,2}$ by adding q1q-1 edges into the color classes arbitrarily, then $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)<\lambda(Y_{n,2,q})$.
 5 Proof of Theorem 2.3
-In this section, we give the proof of Theorem 2.3.
-Let $G$ be an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2,q})$ and $G$ has at most $q\lfloor{n}/{2}\rfloor$ triangles.
-Among such graphs, we choose $G$ as the graph with the maximum spectral radius.
-Our aim is to prove that $G$ is the expected extremal graph $T_{n,2,q}$.
-Note that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
-Then Theorem 2.2 implies that $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles.
+In this section, we are ready to prove Theorem 2.3.
+Equivalently, Theorem 2.3 states that if $G$ is an $n$-vertex graph with at most $q\lfloor{n}/{2}\rfloor$ triangles, then $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$.
+Without loss of generality, we may assume that $G$ attains the maximum spectral radius.
+Our goal is to prove that $G=T_{n,2,q}$.
+The maximality of $G$ implies $\lambda(G)>\lambda(Y_{n,2,q})$.
+So Theorem 2.2 is applicable, and $G$ has exactly $q\lfloor{n}/{2}\rfloor$ triangles.
 Similarly, the previous Lemma 4.1 – Lemma 4.6 still hold in this setting.
-By Lemma 4.7, it follows that $e(G)=\lfloor n^{2}/4\rfloor+q$.
-By Lemma 4.2, we have $e(S)+e(T)=q$ and $e(S,T)=\lfloor n^{2}/4\rfloor$.
-Consequently, we know that $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles.
-Moreover, $G$ is obtained from the bipartite Turán graph $T_{n,2}$ by embedding exactly $q$ edges into the color classes, and there is no triangle among these $q$ edges.
-Combining these facts, it remains to show that the $q$ edges induce a star in the vertex part of size $\lceil{n}/{2}\rceil$.
+By Lemma 4.7, it follows that $e(G)=\lfloor n^{2}/4\rfloor+q$, and $G$ is obtained from $T_{n,2}$ by embedding exactly $q$ edges into the color classes, and there is no triangle among these $q$ edges.
+Combining these facts, it remains to show that these $q$ edges induce a star in the vertex part of size $\lceil{n}/{2}\rceil$.
 Observe that $G$ does not contain $q+1$ pairwise edge-disjoint triangles.
-In particular, if $n$ is even and $q\neq 3$, then $|S|=|T|={n}/{2}$ and setting $k:=q+1$ in [29, Theorem 7], we know that $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$.
-However, the proof [29] follows the line of that of forbidding friendship graph in [9], where the triangle removal lemma is used so that it requires that $q$ is fixed and $n$ is sufficiently large.
+In particular, if $n$ is even and $q\neq 3$, then |S|=|T|=n/2|S|=|T|={n}/{2} and setting $k:=q+1$ in [31, Theorem 7], we know that $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$.
+However, their proof in [31] follows the line of that of forbidding friendship graph in [9], where the triangle removal lemma is used so it requires that $q$ is fixed and $n$ is sufficiently large.
 Actually, the result seems ambiguous whenever $q$ is large (say, when $q\geq\log n$).
@@
 From the above discussion, one of the main difficulties of proving Theorem 2.3 lies in giving a unique characterization of the spectral extremal graph.
-In the sequel, we shall give a complete proof for every $n\geq 300q^{2}$, instead of for $n$ being large enough.
+In the sequel, we shall give a complete proof for every $n\geq 121q^{2}$, instead of for sufficiently large $n$.
 The key ingredient relies on the double-eigenvector technique, which was initially used by Rowlinson [53].
-As far as we know, this technique turns out to be effective for treating spectral extremal graph problems; see [58, 29, 19] and references therein.
+As far as we know, this technique turns out to be effective for treating spectral extremal graph problems; see [61, 31, 19, 35, 66] and references therein.
 The idea in our proof has its root in [31, 19].
 To begin with, we need to introduce two lemmas.
-The first lemma provides an operation of a connected graph that increases the adjacency spectral radius strictly.
+The first lemma provides an operation of a connected graph that strictly increases the adjacency spectral radius.
 Lemma 5.1 (See [58]).
 Assume that $G$ is a connected graph with $u,v\in V$ and $w_{1},\ldots,w_{s}\in N(v)\setminus N(u)$.
-Let $\mathbf{x}=(x_{1},\ldots,x_{n})^{\top}$ be the Perron vector with $x_{v}$ corresponding to the vertex $v$.
+Let $\bm{x}=(x_{1},\ldots,x_{n})^{\mathrm{T}}$ be the Perron vector with $x_{v}$ corresponding to the vertex $v$.
 Let $G^{\prime}=G-\{vw_{i}:1\leq i\leq s\}+\{uw_{i}:1\leq i\leq s\}$.
 If ${x}_{u}\geq{x}_{v}$, then $\lambda(G^{\prime})>\lambda(G)$.
-From the above lemma, we can get the following corollary.
-Corollary 5.2.
-For any $q\geq 2$, we have $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
-The second lemma gives an upper bound on the sum of squares of degrees in a graph with given size. Here, we shall provide an alternative short proof.
+The second lemma gives an upper bound on the sum of squares of degrees in a graph with given size. Here, we provide a two-line simple proof for convenience.
 Lemma 5.2 (See [14, 49]).
@@
 We denote $\lambda=\lambda(G)$ and $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$.
-Let $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})^{\top}$ be the Perron vector of $G^{\prime}$.
+Let $\bm{y}=(y_{1},y_{2},\ldots,y_{n})^{\mathrm{T}}$ be the Perron vector of $G^{\prime}$.
 By the symmetry, we have $y_{v}=y_{w_{1}}$ for every vertex $v\in V(K_{1,a+b})\backslash\{v^{*}\}$.
@@
 Using the double-eigenvector technique, we have
-$\mathbf{x}^{\top}\mathbf{y}\left(\lambda^{\prime}-\lambda\right)$ $=$ $\mathbf{x}^{\top}\left(A(G^{\prime})-A(G)\right)\mathbf{y}$ (2)
+$\bm{x}^{\mathrm{T}}\bm{y}\left(\lambda^{\prime}-\lambda\right)$ $=$ $\bm{x}^{\mathrm{T}}\left(A(G^{\prime})-A(G)\right)\bm{y}$ (2)
 $=$ $\sum_{\{u,v\}\in E(G^{\prime})}(x_{u}y_{v}+x_{v}y_{u})-\sum_{\{u,v\}\in E(G)}(x_{u}y_{v}+x_{v}y_{u})$
@@
 Next, we are going to prove
-$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0.$
+$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0.$
 Since $N_{G}(v^{*})=T\cup N_{H_{1}}(v^{*})$, we have
@@
 $=$ $b(2a+b-\tfrac{3}{2})x_{w_{1}}y_{w_{1}}-b(b+3)x_{v^{*}}y_{w_{1}}.$
-To obtain a contradiction, we would like to show that $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$.
+To obtain a contradiction, we would like to show that $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$.
 It is sufficient to prove that $\frac{x_{v^{*}}}{x_{w_{1}}}<\frac{2a+b-{3}/{2}}{b+3}$.
@@
 $\frac{x_{v^{*}}}{x_{w_{1}}}\leq\frac{\lambda}{\lambda-a}<1+\frac{1}{b+3}<\frac{2a+b-{3}/{2}}{b+3},$
-which leads to $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$ and $\lambda^{\prime}>\lambda$, a contradiction. Thus, $H_{1}$ is a star. ∎
-In Lemma 5.4, we expect to prove that $H_{i}$ is a star, instead of a $4$-cycle.
-But the above argument seems invalid for $H_{1}\cong C_{4}$, so we need to make an easy comparison between these two cases.
-If $n$ is odd, then $|S|={(n-1)}/{2}$ and $|T|={(n+1)}/{2}$.
-Since $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles, we know that the $q$ edges must be embedded into the larger part $T$.
-By Lemma 5.4, we must embed a star $K_{1,q}$ or a cycle $C_{4}$ (in the case of $q=4$) into $T$.
-By direct computation, one can check that embedding $K_{1,4}$ to $T_{n,2}$ attains the maximum spectral radius.
+which leads to $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$ and $\lambda^{\prime}>\lambda$, a contradiction. Thus, $H_{1}$ is a star. ∎
+In Lemma 5.3, we expect to prove that $H_{i}$ is a star, instead of a $4$-cycle. So we need to make an easy comparison between these two cases.
+If $n$ is odd, then $|S|=\frac{n-1}{2}$ and $|T|=\frac{n+1}{2}$.
+Since $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles, the $q$ edges must be embedded into $T$.
+By Lemma 5.3, we must embed a star $K_{1,q}$ or a cycle $C_{4}$ (in the case of $q=4$).
+By computation, embedding a copy of $K_{1,4}$ attains the maximum spectral radius.
 Indeed, we can compute that $\lambda(T_{n,2,4})$ is the largest root of
@@
 It is not hard to check that $\lambda(G)<\lambda(T_{n,2,4})$. This completes the proof for odd $n$.
-If $n$ is even, then $|S|=|T|={n}/{2}$.
+If $n$ is even, then |S|=|T|=n/2|S|=|T|={n}/{2}.
 Similarly, the embedded graphs $H_{1}$ and $H_{2}$ are stars or $4$-cycles by Lemma 5.3.
@@
 Let $n$ be even and $G^{\prime}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the part $S$.
-Let $\lambda^{\prime}=\lambda(G^{\prime})$ and $\mathbf{y}$ be the Perron vector of $G^{\prime}$.
+Let $\lambda^{\prime}=\lambda(G^{\prime})$ and $\bm{y}$ be the Perron vector of $G^{\prime}$.
 Then
@@
 Without loss of generality, we may assume that $x_{T}\geq x_{S}$.
-Let $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$ and $\mathbf{y}=\left(y_{1},y_{2},\ldots,y_{n}\right)^{\top}$ be the Perron vector of $G^{\prime}$.
+Let $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$ and $\bm{y}=\left(y_{1},y_{2},\ldots,y_{n}\right)^{\mathrm{T}}$ be the Perron vector of $G^{\prime}$.
 To show $\lambda^{\prime}>\lambda$, we shall prove
-$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0.$
+$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0.$
 Since $G$ is the graph with the maximum spectral radius, we have $\lambda\geq\lambda^{\prime}>\frac{n}{2}$.
@@
 Considering the above inequalities, we can obtain
-$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
+$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
 $=$ $\sum_{\{u,v\}\in E(H)}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)-\sum_{\{u,v\}\in E\left(H_{1}\right)\cup E\left(H_{2}\right)}\lambda\lambda^{\prime}\left(x_{u}y_{v}+x_{v}y_{u}\right)$
@@
 By computation, we can check that $q_{1}q_{2}\geq q-1$ and
-$c_{1}-c_{2}=\frac{2q_{1}q_{2}\lambda-q^{3}-q^{2}}{\lambda(\lambda-q)}>\frac{(q-1)n-q^{3}-q^{2}}{(n/2+q)n/2}>\frac{1}{10n},$
+$c_{1}-c_{2}=\frac{2q_{1}q_{2}\lambda-q^{3}-q^{2}}{\lambda(\lambda-q)}>\frac{(q-1)n-q^{3}-q^{2}}{(n/2+q)n/2}>\frac{1}{5n},$
 where we used the assumption ${n}/{2}<\lambda<{n}/{2}+q$ and $n\geq 121q^{2}$.
@@
 Combining with the assumption that $x_{T}\geq x_{S}$, we get
-$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
+$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
 $\geq$ $c_{1}x_{T}y_{T}-c_{2}x_{S}y_{S}$
-$\geq$ $\left(c_{2}+\frac{1}{10n}\right)x_{S}\left(y_{S}-\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}\right)-c_{2}x_{S}y_{S}$
-$>$ $\frac{x_{S}}{10n}\cdot\frac{\lambda^{\prime}y_{V}}{n+q}-(2q-1)x_{S}\cdot\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V},$
+$\geq$ $\left(c_{2}+\frac{1}{5n}\right)x_{S}\left(y_{S}-\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}\right)-c_{2}x_{S}y_{S}$
+$>$ $\frac{x_{S}}{5n}\cdot\frac{\lambda^{\prime}y_{V}}{n+q}-(2q-1)x_{S}\cdot\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V},$
 where the last inequality holds by $y_{S}=\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}>\frac{\lambda^{\prime}y_{V}}{n+q}$ and $c_{2}<2(q-1)+\frac{q^{2}}{n/2-q}$.
-Hence, we can obtain $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0$ for $n\geq 300q^{2}$.
+Hence, we can obtain $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0$ for $n\geq 121q^{2}$.
 This a contradiction with $\lambda^{\prime}>\lambda$.
@@
 Proof of Theorem 2.5.
-Assume that $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, where $s\in\mathbb{N}$.
-If $s=0$, then there is nothing to show.
+Assume that $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, where $s\geq 1$.
 If $s=1$, then $G\neq T_{n,2}$.
-Theorem 1.3 implies $G$ has $\lfloor n/2\rfloor-1$ triangles.
-For the case $s=2$, it was proved in [27] that $G$ contains at least $n-3$ triangles.
+Theorem 1.3 implies $G$ has $\lfloor\frac{n}{2}\rfloor-1$ triangles.
+For $s=2$, the authors [29] proved that $G$ has at least n3n-3 triangles.
 Now, we consider the case $s\geq 3$.
-We may assume that $t(G)\leq s{(n-1)}/{2}$.
+We may assume that $t(G)\leq\frac{1}{2}s{(n-1)}$.
 Otherwise, we are done.
 Claim 1.
-We have $e(G)\geq\lfloor n^{2}/4\rfloor-3s$.
-There is a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6s$, $e(S,T)>\lfloor n^{2}/4\rfloor-9s$ and ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$.
+We have $e(G)>\lfloor n^{2}/4\rfloor-3s$.
+There is a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6s$, $e(S,T)>\lfloor n^{2}/4\rfloor-9s$ and $\frac{n}{2}-3\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+3\sqrt{s}$.
 Proof of Claim 1.
-Since $\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>{(n-1)}/{2}$, by Lemma 3.2, we have
-$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-3s.$
+Since $\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>\frac{1}{2}(n-1)$, by Lemma 3.4, we have
+$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-3s.$
 If $G$ is $6s$-far from being bipartite, then using Theorem 3.5, we get
@@
 $e(S)+e(T)<6s.$
-Consequently, we get $e(S,T)=e(G)-e(S)-e(T)>\lfloor{n^{2}}/{4}\rfloor-9s$.
+Consequently, we get
+$e(S,T)=e(G)-e(S)-e(T)>\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-9s.$
 By the symmetry, we may assume that $|S|\leq|T|$.
-Suppose on the contrary that $|S|<{n}/{2}-3\sqrt{s}$.
-Then $|T|=n-|S|>{n}/{2}+3\sqrt{s}$.
+Suppose on the contrary that $|S|<\frac{n}{2}-3\sqrt{s}$.
+Then $|T|=n-|S|>\frac{n}{2}+3\sqrt{s}$.
 It follows that $e(S,T)\leq|S||T|\leq\lfloor n^{2}/4\rfloor-9s$, which is a contradiction.
-Thus, we have ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$.
+Thus, we have $\frac{n}{2}-3\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+3\sqrt{s}$.
 ∎
@@
 Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$.
-If $e(S)+e(T)\leq s-1$, then all triangles of $G$ can be covered by a vertex set of at most $s-1$ vertices, contradicting with $\tau_{3}(G)\geq s$.
-Now, suppose that $e(S)+e(T)\geq s+1$.
-Note that each edge of $G[S]$ or $G[T]$ can yield at least ${n}/{2}-3\sqrt{s}$ triangles in $K_{|S|,|T|}$.
-From the above discussion, we know that $e(S)+e(T)<6s$ and $|S||T|-e(S,T)\leq\lfloor n^{2}/4\rfloor-e(S,T)<9s$.
-Then $G[S,T]$ misses less than $9s$ edges between $S$ and $T$, and each missing edge destroys at most $6s$ triangles.
-So $G$ contains more than $(s+1)({n}/{2}-3\sqrt{s})-9s\cdot 6s>(s+1){n}/{2}-56.5s^{2}\geq s{n}/{2}$ triangles, where the last inequality holds for every $n\geq 113s^{2}$.
-This contradicts with the assumption $t(G)\leq s{(n-1)}/{2}$.
+If $e(S)+e(T)\leq s-1$, then all triangles of $G$ can be covered by a vertex set of at most s1s-1 vertices, contradicting with $\tau_{3}(G)\geq s$.
+Now, suppose that $e(S)+e(T)=k$ for some integer $k\geq s+1$.
+Note that each edge of $G[S]$ or $G[T]$ can yield at least ${n}/{2}-3\sqrt{s}$ triangles.
+From Claim 1, we know that $|S||T|-e(S,T)\leq\lfloor n^{2}/4\rfloor-e(S,T)<9s$.
+Then $G[S,T]$ misses at most (9s1)(9s-1) edges between $S$ and $T$, and each missing edge destroys at most $k$ triangles.
+So we get $t(G)\geq k(\frac{n}{2}-3\sqrt{s})-(9s-1)k\geq k(\frac{n}{2}-3\sqrt{s}-9s+1)>\frac{1}{2}s(n-1)$, where the last inequality holds for every $s\geq 3$ and $n\geq 28s^{2}$.
+This contradicts with the assumption $t(G)\leq\frac{1}{2}s(n-1)$.
 Thus, we have $e(S)+e(T)=s$, as claimed.
 ∎
-From the above claim, we obtain $e(S,T)=e(G)-s>\lfloor n^{2}/4\rfloor-4s$. Then
+From the above claims, we obtain $e(S,T)=e(G)-s>\lfloor n^{2}/4\rfloor-4s$. Then
 $\frac{n}{2}-2\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+2\sqrt{s}.$
-Specifically, if $s=3$, then we have $\lceil n/2\rceil-3\leq|S|,|T|\leq\lfloor n/2\rfloor+3$.
-Since $e(S)+e(T)=s$ and there are less than $4s$ missing edges in $G[S,T]$, we know that $G$ contains at least $s(\lceil n/2\rceil-3)-4s^{2}\geq sn/2-5s^{2}$ triangles.
-If $s\geq 4$, then a similar argument shows that $G$ has at least $s({n}/{2}-2\sqrt{s})-4s^{2}\geq s{n}/{2}-5s^{2}$ triangles.
+If $s=3$, then $\lceil\frac{n}{2}\rceil-3\leq|S|,|T|\leq\lfloor\frac{n}{2}\rfloor+3$.
+Since there are less than $4s$ missing edges in $G[S,T]$, Claim 2 implies that $t(G)\geq s(\lceil\frac{n}{2}\rceil-3)-4s\cdot s\geq\frac{1}{2}sn-5s^{2}$.
+If $s\geq 4$, then a similar argument shows that $t(G)\geq s(\frac{n}{2}-2\sqrt{s})-4s^{2}\geq\frac{1}{2}sn-5s^{2}$, as needed.
 ∎
@@
 Problem 7.1.
-Finding the limit superior of $C$ such that
-$t(G)\geq Cn^{2}\left(\lambda-\frac{n}{2}\right).$
+What is the best constant $C>0$ such that $t(G)\geq Cn^{2}\left(\lambda-\frac{n}{2}\right)$?
 Remark.
 By the clique density theorem of Reiher [52], we know that if $\varepsilon>0$ and $G$ is a graph on $n$ vertices with $e(G)\geq(\frac{1}{4}+\varepsilon)n^{2}$, then $G$ has at least $(\frac{1}{2}+o(1))\varepsilon n^{3}$ triangles, where $o(1)\to 0$ whenever $\varepsilon\to 0$.
-With aid of this powerful result, one can easily prove that if $G$ has at least $(\frac{1}{4}+\varepsilon)n^{2}$ edges, where we set $\varepsilon=\frac{1}{2}(\frac{\lambda}{n}-\frac{1}{2})$, then $G$ has at least $(\frac{1}{4}+o(1))n^{2}(\lambda-\frac{n}{2})$ triangles.
+With the aid of this powerful result, one can easily prove that if $G$ has at least $(\frac{1}{4}+\varepsilon)n^{2}$ edges, where we set $\varepsilon:=\frac{1}{2}(\frac{\lambda}{n}-\frac{1}{2})$, then $G$ has at least $(\frac{1}{4}+o(1))n^{2}(\lambda-\frac{n}{2})$ triangles.
 Moreover, the graph $T_{n,2,1}$ shows that the coefficient $\frac{1}{4}$ is achievable.
-Very recently, Shengtong Zhang told us that taking $G=T_{n,3}$ yields $C\leq\frac{2}{9}$.
+Recently, Shengtong Zhang told us that taking $G=T_{n,3}$ yields $C\leq\frac{2}{9}$.
 Another question as to how many triangles appear at least in any $n$-vertex graph with given spectral radius turns out to be interesting.
-For instance, for any graph $G$ with $\lambda(G)=\gamma n$, where $\gamma\in(\frac{1}{2},\frac{2}{3}]$, that is, $\lambda(T_{n,2})<\lambda(G)\leq\lambda(T_{n,3})$, there might exist a graph, say $L$, whose shape is approximately between $T_{n,2}$ and $T_{n,3}$, and $\lambda(L)=\gamma n$ such that $t(L)\leq t(G)$.
-Formally, we now define the minimizer $L$ as below.
+For instance, for any graph $G$ with $\lambda(G)=\gamma n$, where $\gamma\in(\frac{1}{2},\frac{2}{3}]$, that is, $\lambda(T_{n,2})<\lambda(G)\leq\lambda(T_{n,3})$, there might exist a graph, say $L$, whose shape is approximately between $T_{n,2}$ and $T_{n,3}$ such that $\lambda(L)=\gamma n$ and $t(L)\leq t(G)$.
+Formally, we now define the minimizer $L$ as follows.
 For $s\in\mathbb{N}$ and $\alpha>0$, we define $L_{n,s,\alpha}$ as a complete multipartite graph with $V(L_{n,s,\alpha})=V_{1}\cup\cdots V_{s}\cup V_{s+1}$, where $|V_{1}|=\cdots=|V_{s}|=\frac{n(1+\alpha)}{s+1}$ and $|V_{s+1}|=\frac{n(1-s\alpha)}{s+1}$.
-Next, we conjecture that $L_{n,s,\alpha}$ contains the minimum number of triangles among all graphs with given spectral radius.
+Next, we speculate that $L_{n,s,\alpha}$ contains the minimum number of triangles among all graphs with given spectral radius.
 Conjecture 7.2 (Spectral triangle density conjecture).
@@
 Nosal [50] proved that every graph $G$ with $m$ edges and $\lambda(G)>\sqrt{m}$ contains a triangle; see, e.g., [41, 47].
-Recently, Ning and Zhai [45] showed that if $\lambda(G)\geq\sqrt{m}$, then $t(G)\geq\lfloor\frac{\sqrt{m}-1}{2}\rfloor$, unless $G$ is a complete bipartite graph.
+In 2023, Ning and Zhai [48] showed that if $\lambda(G)\geq\sqrt{m}$, then $t(G)\geq\lfloor\frac{\sqrt{m}-1}{2}\rfloor$, unless $G$ is a complete bipartite graph.
 We propose the following conjecture.
@@
 Thus, the difficulty of Conjecture 7.3 lies in deriving the tight bound $t(G)\geq\frac{m-1}{2}$.
-Recall in Theorem 1.2 that for every $q<n/2$, the graph $G$ with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles.
-While, in our Theorem 2.2, we proved the spectral version for $n\geq 300q^{2}$.
-Under our framework, it seems possible to improve this bound slightly.
-For the sake of clarity, we do not make any serious attempt to optimize the absolute constant in our statements and proofs.
-It is ambitious to extend our result to the following wide range.
+Theorem 1.2 asserts that every graph $G$ with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles for every $q<n/2$.
+In Theorem 2.2, we proved a spectral version for all $q\leq\frac{1}{11}\sqrt{n}$.
+This bound on $q$ is tight up to a constant factor (Example 3.3).
+Under our framework, it seems feasible to improve the constant factor slightly.
+However, new ideas are needed to determine the exact bound on $q$.
+It is ambitious to propose the following problem.
 Problem 7.4.
-Whether Theorems 2.2 and 2.3 hold for every $q<n/2$? If not, what is the limit superior of $\delta$ such that the spectral results hold for $q<\delta n$?
-In 2017, Pikhurko and Yilma [48, Theorem 1.4] proved that for every color-critical graph $F$ with $\chi(F)=r+1$, there exist $\delta=\delta(F)>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and $q<\delta n$, if $G$ is an $n$-vertex graph with $e(T_{n,r})+q$ edges and it has the minimum number of copies of $F$, then $G$ is obtained from $T_{n,r}$ by adding $q$ new edges.
-Theorem 1.2 gives $\delta(K_{3})=1/2$.
-More generally, a result of Lovász and Simonovits [34, 35] implies that $\delta(K_{r+1})=1/r$ for every $r\geq 2$.
-Moreover, Pikhurko and Yilma [48] also determined the limit superior of $\delta$ for a number of color-critical graphs, such as, odd cycles $\delta(C_{2k+1})=1/2$, and cliques with one edge removed $\delta(K_{r+2}-e)=(r-1)/r^{2}$.
-In a forthcoming paper, we will use a variety of different methods to establish a spectral counting result for any color-critical graph.
-As the techniques are quite complicated, we intend to present it separately $\heartsuit$.
-Declaration of competing interest
-The authors declare that they have no conflicts of interest to this work.
-Data availability
-No data was used for the research described in the article.
+What is the supremum $\delta\in(0,0.8)$ such that Theorem 2.2 holds for every integer $q<\delta\sqrt{n}$? Additionally, what is the optimal range of $q$ in Theorem 2.3?
+In 2017, Pikhurko and Yilma [51] proved that for every color-critical graph $F$ with $\chi(F)=r+1$, there exist $\delta=\delta(F)>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and $q<\delta n$, if $G$ is an $n$-vertex graph with $e(T_{n,r})+q$ edges and it achieves the minimum number of copies of $F$, then $G$ is obtained from the Turán graph $T_{n,r}$ by adding $q$ new edges.
+Furthermore, by Theorem 1.2, we see that $\delta(K_{3})=\frac{1}{2}$.
+A result of Lovász and Simonovits [37, 38] implies that $\delta(K_{r+1})=\frac{1}{r}$ for every $r\geq 2$.
+Moreover, Pikhurko and Yilma [51] also determined the exact value of $\delta$ for a number of color-critical graphs, such as, odd cycles $\delta(C_{2k+1})=\frac{1}{2}$, and cliques with one edge removed $\delta(K_{r+2}-e)=\frac{r-1}{r^{2}}$.
+Similarly, it is also meaningful to determine the exact value of $\delta(F)$ for various graphs $F$ in the spectral setting.
 Acknowledgements
-The authors are thankful to Xiaocong He and Loujun Yu for carefully reading an early manuscript of this paper.
-Yongtao Li would like to thank Dheer Noal Desai for the inspiring discussions on spectral graph problems over the years.
-Li was supported by the Postdoctoral Fellowship Program of CPSF (No.
-GZC20233196), Feng was supported by the NSFC grant (No.
-12271527), and Peng was supported by the NSFC grant (No.
-11931002).
+We are grateful to Xiaocong He and Loujun Yu for carefully reading an early draft of the paper.
+Yongtao Li would like to thank Dheer Noal Desai for inspiring discussions on spectral graph problems.
+Lihua Feng was supported by the National Natural Science Foundation of China (Nos.
+12271527 and 12471022), and Yuejian Peng was supported by the National Natural Science Foundation of Hunan Province (No.
+2025JJ30003) and the National Natural Science Foundation of China (No.
+12571363).